### v1

# Formalising the intentional stance: attributing goals and beliefs to stochastic processes

Abstract
We are concerned with the behaviour of stochastic systems with inputs and outputs, and how this might relate to the pursuit of a goal. We model this using what we term transducers, which are a mathematical object that captures only the external behaviour of such a system and not its internal state. We present a framework for reasoning about the optimality of such a process, when it is coupled to a ‘teleo-environment’ consisting of another transducer that also embodies a success criterion. We find that (globally) optimal transducers have a property closely related to Bellman’s theorem: a transducer that is optimal in one time step will again be optimal in the next time step, but with respect to a different environment (obtained from the original one by a modified version of Bayesian filtering). We also consider bounded rationality and its relationship to constrained optimality, which in our framework means optimal within some subset of all transducers. We describe a condition that is sufficient for such a subset to have this Bellman property. Additionally, we show that a policy is deterministic if and only if there exists a teleo-envionment for which it is uniquely optimal among the set of all transducers; this is at least conceptually related to classical representation theorems from decision theory. This need not hold for constrained subsets; we give an example of this related to the so-called absent-minded driver problem. All of the formalism is defined using coinduction, following the style proposed by Czajka [9].
Keywords: mathematical formulations of agency, controlled stochastic processes, as-if agency, intentional stance, POMDPs, applications of coinduction
Classification: 91E99, 93-10, 93E03
1 Introduction
We are interested in the question of what does it mean for a system to be an agent? More specifically, we are interested in the relationship between what we might call mechanistic descriptions of a system (explanations of its behaviour in terms of its internal state) and what we call normative-epistemic descriptions. These are descriptions in terms of a combination of norms and beliefs. For example, we might ascribe a goal to the system and beliefs about its surroundings, and explain its behaviour in terms of those. Both types of description concern the externally observable behaviour of a system, but they are quite different in nature, and we want to understand where this difference comes from and exactly how they relate. We take a position similar to [11] in that we regard in intentional descriptions as a stance, i.e. an optional perspective that can be taken by an observer, rather than an inherent property of a system. Our goal is to express this idea mathematically.
In order to do this, we propose a simple mathematical framework in which systems can be described in both mechanistic and normative-epistemic terms, and the relationships between the two can be understood. We don’t aim for the most general framework possible but for the simplest one in which our main points can be expressed. We express this framework in the mathematical language of coinduction, and indeed, a secondary purpose of this paper is to demonstrate the application of coinduction to stochastic systems.
In the remainder of this section we give an overview of our framework and a description of our results, as well as a brief explanation of our coinductive approach. In Section 2 we describe relations to previous approaches, including Dennett’s work, representation theorems for expected utility theory, the free energy principle, computational mechanics and inverse reinforcement learning, as well as our own work on the ‘as-if’ approach to cognition. The mathematical content of the paper begins in Section 3.
1.1 Outline of our approach and results
In this section we give a somewhat detailed summary of the paper’s results, using a minimum of technical language.
Consider the following situation: A person who we will refer to as ‘the theorist’ is studying a physical system. Although we use the term ‘physical system’, a suitable image to have in mind might be that of an animal, or even better, a robot. (We can imagine that the robot was designed by someone other than the theorist, for a purpose that the theorist doesn’t necessarily know about.) The theorist’s goal is to understand whether the system they are dealing with is an agent. Or, more accurately, whether it can be consistently interpreted as if it were an agent, in the sense that it takes actions that are consistent with trying to achieve some particular goal. There might be more than one way in which the system can be so interpreted; if so, the theorist is interested in all of them.
In order to reason about physical systems in such a situation, we make the following assumptions: the system has a set of possible outputs or ‘actions’ $A$ and a set of possible inputs or ‘sensor values’ $S$ (both finite). Time proceeds in discrete steps and the system’s behaviour may be stochastic. We can model its externally observable behaviour in terms of a controlled stochastic process, that is, a conditional probability distribution over infinite sequences of outputs or actions $a_{1},a_{2},\dots$ given infinite sequences of inputs or sensor values $s_{1},s_{2},\dots$. This conditional distribution must have the property that the output at time $t$ is independent of all inputs received after time $t-1$. (See [20] for a similar approach.) We refer to this as the causality condition. We will see later on that controlled stochastic processes can be defined in a different way using the mathematical framework of coinduction, and we will use coinductive reasoning throughout the paper. In the context of our coinductive definition we refer to controlled stochastic processes as transducers.
(a) mechanistic description
(b) normative-epistemic description
Figure 1: Two ways in which a theorist might describe the behaviour of a physical system. $(a)$ in a mechanistic description the theorist posits an internal state to the system such that its dynamics can explain the observable behaviour; in the case of a robot this could include hypothesised workings of internal control systems in addition to the literal mechanical gears depicted here. We formalise mechanistic descriptions in terms of stochastic Moore machines in Section 3.3.4. $(b)$ in a normative-epistemic description the theorist posits an environment that the system believes itself to be in (according to the theorist’s interpretation) and a goal that it seeks to achieve — in this case the apple — such that the system’s behaviour can be explained by acting optimally in order to reach the goal. Normative-epistemic descriptions are the focus of our paper and are formalised in terms of teleo-envionments in Section 4. We emphasise that these are two different kinds of explanation for the same behaviour of the same system, and not two different types of system or two competing hypotheses.
We assume that the theorist has access to a controlled stochastic process representing their system’s observable behaviour. This means that the theorist is not limited to observing a single sample of the system’s behaviour given a single sequence of inputs; instead they know111or at least, have a model of how the system would behave given any number of different input sequences, over any number of trials.
Given this knowledge, one question one could ask is, “given the known externally observable behaviour, what kinds of physical mechanism might exist inside the system that would give rise to its behaviour?” Although we will briefly discuss this question in Section 3.3.4 it will not be our main focus.
Instead, we are concerned with the question “is the system’s behaviour consistent with taking optimal actions in order to achieve some goal?” The relationship between these two questions is illustrated in Figure 1.
To address this, we introduce the notion of teleo-environment. This is also a transducer, but with the opposite interface to the system: where the system takes input from a set $S$ and gives outputs in a set $A$, its environment takes actions $A$ as inputs and gives sensor values $S$ as outputs, so that the two systems can be coupled. (They are assumed to couple in such a way that they give their respective outputs simultaneously, as opposed to alternating.)
In addition to this, a teleo-environment comes equipped with a notion of success. The idea is that success is an event that may or may not occur on any given time step, and the agent’s goal (as attributed by the theorist) is to maximise the probability that success occurs at least once. The agent does not necessarily know whether success has been achieved or not on any given time step, and hence can be ‘motivated’ to continue ‘trying’ to succeed even if success has already occurred. This differs somewhat from the usual formulation of partially observable Markov decision processes (POMDPs) in terms of exponentially discounted reward, but is arguably more intuitive and results in an arguably simpler mathematical framework; it also has some interesting consequences. Hence we define a teleo-environment as a transducer from $A$ to $S\times\{\bot,\top\}$ where $\bot$ is a ‘lack of success’ signal, and $\top$ is a ‘success’ signal. This notional ‘telos’ channel is not directly observable by the system, but forms part of our concept of optimality.
It is worth noting that our notion of ‘goal’ is inherently tied to a notion of ‘environment’, since teleo-environments embody both in a way that can’t easily be separated. When we vary the goal in order to ask about the set of all goals for which a given agent is optimal, we vary the environment as well. (This is in contrast to the usual situation in inverse reinforcement learning, in which the environment and goal are separate, and the environment is known but the goal is not.)
Our position is that, intuitively, a teleo-environment can be taken to represent a particular normative-epistemic state, meaning a description of a system’s behaviour in terms of goals and beliefs. We regard it as describing both what the agent believes about its environment (as attributed to it by the theorist), and what it is (interpreted as) trying to do. The idea is that if a policy $\pi$ is optimal for a teleo-environment $\varepsilon$, it’s consistent to claim that $\pi$ behaves as it were an ideally rational agent, with beliefs and goal represented by $\varepsilon$. In this sense, we can treat $\varepsilon$ as a permissible ‘mental state attribution’ for $\pi$. The teleo-environment $\varepsilon$ does not necessarily represent the dynamics of the true environment, but rather the environment that the transducer $\pi$ ‘believes’ it is in, according to the normative-epistemic state attributed to it by the theorist.
Since $\varepsilon$ is not unique, multiple distinct mental state attributions may be applied to the same system. We see this as a strength of our approach rather than a weakness, since it captures the idea that not only is taking the intentional stance a choice, but that there can be choices in how it is applied as well.
Within this simplified model of ideal rationality, we can ask: as a system ‘evolves’ its policy through interaction with its environment, how do the permissible mental state attributions change? By ‘evolving’ we mean the following: a transducer, i.e. a controlled stochastic process, can be seen as a conditional distribution of output sequences given input sequences, subject to a causality condition. However, once the transducer has emitted an output and received an input we can obtain a new distribution over output sequences given input sequences, by conditioning on the input-output pair and shifting both sequences by one time step, as described in sections 3.1 and 3.2.
We show in section 5 that, if $\pi$ is optimal for $\varepsilon$, then after ‘evolution’ with a sensorimotor pair $(s,a)$ the evolved policy $\pi^{\prime}$ is optimal for a similar evolution $\varepsilon^{\prime}$ of $\varepsilon$. We argue in that section that this can be seen as a form of Bayesian filtering. In other words, among the many beliefs that are permissibly attributable to the system after it receives new information, one is guaranteed to look like an update of the previous attributed belief according to Bayes’ rule. (There is some subtlety about exactly what data is conditioned on, however, which will be discussed shortly.)
We propose a general intuition along the following lines. Consider any notion of optimality for a control policy $x$ in an environment $y$, and some notion of updating over time for the policy and environment (we will be deliberately vague about what this ‘updating’ means in general). Suppose something like Bellman’s principle holds, namely that if a policy $x$ is optimal for $y$, then the evolution $x^{\prime}$ by a single time step will be optimal for the evolution $y^{\prime}$ of $y$. Again, $y$ will function as a permissible ‘mental state’ attribution for $x$ (in the sense that no inconsistency arises from supposing that $x$ is an ideal agent trying to solve the problem represented by $y$) and similarly for $y^{\prime}$ and $x^{\prime}$. Consequently, the ‘update’ operator that we apply to $y$ to obtain $y^{\prime}$ can be seen as updating beliefs attributed to $x$, to obtain beliefs that can be attributed to $x^{\prime}$, given any information recieved by $x$ during the update. In this sense the ‘update’ rule for $y$ can be seen as a model of idealised rationality, which generalises Bayes’ rule in some sense.
When we work through the details for our teleo-environment model, we obtain an interesting insight. One might reasonably guess that a Bayesian model of rational mental state update might emerge from our model. It does, but with an important nuance: so to speak, the system is legitimated to update its ‘belief’ in a Bayesian manner, conditioning on the sensorimotor signals observed in the previous time step, provided that it additionally conditions on not having achieved its goal in the previous time step. That is, it conditions on its reieved sensor value $s\in S$, but it also conditions on $\bot$, the lack-of-success signal, even though it doesn’t actually have access to the success signal, and indeed the true value of the success signal might have been $\top$. We call this ‘value-laden filtering’, and it arises from the relevant Bellman-like principle (in Theorem 5.0.2).
This is a little surprising, since an external theorist, observing the sensorimotor interactions between the system and its environment (without any additional information about whether a goal had been achieved), and applying Bayesian reasoning themselves to predict future interactions, would condition only on the sensorimotor signals. In this way, the properties of the optimality criterion in our model mean that we can derive a permissible rule for pragmatically optimal ‘mental state’ change (namely, value-laden filtering) which differs subtly from the standard Bayesian picture, whilst still being recognisable as Bayesian filtering. In future, it may be of interest to explore further how different pragmatic notions of optimality can induce different permissible update rules for mental state, and how these rules relate to ‘purely normative’ models of ideal mental state update.
Another question one might want to ask is whether a given behaviour can be specified by a given teleo-environment. That is, for a system with given behaviour, does there exist a teleo-environment for which it is uniquely optimal? We explore this question in Section 6. Our results here are as one might expect from a framework based on utility-maximisation. Firstly, a behaviour can only be uniquely optimal for a given teleo-environment if it is deterministic, since otherwise it would be a mixture of two or more deterministic behaviours, each of which would also be optimal. Secondly, the specifiable behaviours are exactly the deterministic ones: if the behaviour is deterministic then there exists a teleo-environment for which it is unqiuely optimal. (There may be many such teleo-environments, but in particular there is always one in which the system must perform exactly the specified behaviour, otherwise it immediately permanently loses all hope of achieving success.) This is shown in Theorem 6.3.2. Thus, at least when it comes to deterministic behaviour, there is no difference in the set of behaviours that can be specified in a normative-epistemic way and the set of behaviours that can be specified by a framework based on internal state.
This is in a sense a ‘no-go theorem’. One might initially have the intuition that systems that are ‘agents’ (in the sense of being optimal for some teleo-environment) might have special features to their behaviour, which would then allow one to test the hypothesis that some system is an agent by observing its behaviour. (See [31], which takes a Bayesian approach along these lines.) Although Theorem 6.3.2 is not a mathematically surprising result it does suggest a challenge that such an approach would need to overcome.
Finally, we touch on the issue of bounded rationality. A transducer may fail to be be optimal for a given teleo-environment $\varepsilon$ while still being optimal for $\varepsilon$ within some constrained class of transducers, meaning simply a subset of all transducers. These constrained classes of transducers are intended to model limitations in the cognitive capacity of a type of agent, such as a limited supply of memory, or of randomness. For some kinds of limitation (i.e. some constrained classes of transducers), the value-laden Bellman property still holds. (In particular, it holds for constrained classes obeying the so-called exchange property, Definition 3.3.1.) For agents that are optimal within such a class it is permissible to attribute normative-epistemic states that update by value-laden filtering, as described above. However, this is not true for all constrained classes of interest. We demonstrate this by applying our framework to a version of the ‘absent-minded driver,’ [32] a classic problem from decision theory that involves making decisions without memory. The optimal solution to the absent-minded driver problem is stochastic rather than deterministic, and as a transducer it is not possible to give it a mental state attribution that updates by value-laden filtering.
1.2 Formal methodology: the coinductive approach
Our formal treatment uses coinductive reasoning to construct the objects we call transducers. While the coinductive formalism might not be familiar to most readers, it corresponds closely to the intuition of ‘moving step-by-step’ and produces elegant proofs. For any readers wishing to become familiar with the formal details of coinductive reasoning we recommend [9] for a set-theoretic justification for the approach we use in this paper, which treats coinduction as a first-class primitive. Alternatively, [22] provides a comprehensive introduction from the perspective of category theory, and [1] a type-theoretic approach.
Intuitively, while induction lets us define and reason about objects with a finite internal structure, such as natural numbers or finite trees, coinduction lets us define and reason in a similar way about objects with an infinite internal structure. In practice this amounts to accepting objects that do not start at an inductive base case, such as trees without leaf nodes, which are necessarily infinite. Coinduction is the category-theoretic dual of mathematical induction, and therefore many of the concepts associated with induction have coinductive analogs. The main mathematical consequence of this is that one can define functions into coinductive objects using the function being defined [9, Theorem 2.4]; [1, Theorem 7] and prove theorems about relations that are coinductively defined [9, Theorem 4.29] [^2] No theorem for this is needed in [1] due to type theory treating propositions as types. or equality [9, Theorem 4.33][1, Theorem 18] using coinductive hypotheses (by analogy to inductive hypotheses).
Note that we could have derived the same results without using coinduction, but we chose to use the coinductive formalism in order to emphasise the dynamical relations between transducers, permitting an elegant set of definitions and proofs.
2 Relation To Previous Work
The current work is part of a research programme of ‘as-if’ agency. The idea of the ‘as-if’ approach [26, 27] is to formally address what it means to treat a system as if it were an agent, while remaining agnostic about whether or not there are additional criteria that must be met for a system to be deemed a ‘true’ agent.
This idea can be seen as an attempt to formalise Dennett’s intentional stance [10, 12, 13]; in doing so, we are also able to formally model relationships between (analogues of) Dennett’s intentional stance and his physical stance. In the current work, the physical stance corresponds roughly to our treatment of stochastic Moore machines and unifilar machines (Section 3.3), while the intentional stance correpsonds to our exploration of the relationship between transducers and teleo-environment, explored in Section 4 and throughout the paper.
We emphasise, though, that while our mathematical treatment has some features that are strongly suggestive of Dennett’s ideas, and is inspired by Dennett’s work, we don’t consider it a full formalisation of Dennett’s work as such, and it isn’t intended to address everything Dennett is concerned with.
In the past there have been significant disagreement about what sorts of system should be regarded as ‘agents’. Some authors (e.g. [34, p.34] or [43]) categorise even simple feedback systems (such as thermostats or guided torpedoes) as agents. Others (e.g. [14] or [23]) explicitly reject such a characterisation, claiming that thermostats may behave (in some ways) as if they were agents, but are not truly agents. A discussion of this general debate can be found in [25]. Currently, echoes of this disagreement can be seen in disputes about the extent to which AI systems such as large language models (LLMs) should be understood as agents. The idea of ‘as-if’ agency is that we should not ignore simple systems such as thermostats when studying agency, even if the skeptics are correct and thermostats lack some important property. It is still valuable to have a theory that explains in what sense even a thermostat behaves as though it were an agent.
The ‘as-if’ approach involves considering the match between a system’s physical behaviour, and the prescriptions of a normative theory of ‘ideal’ rationality [26, 27]. As outlined in [27], “[I]f the causal effects exerted by some system $X$ on its environment are roughly rational decisions under some formal theory of rationality $R$ parameterised by some cognitive parameters $C$, we will say that $X$ is an as-if agent with respect to $(R,C)$.”
We will not define what a ‘formal theory of rationality’ is, but our notion of teleo-environment (Section 4) will serve as an example.
The theory we describe in this paper is not utility theory (as suggested in [27]), but instead a simplified normative model of goal-belief rationality, in which the agent is treated as if it is trying to maximise the probability with which a goal event will occur. Some philosophical motivation for this is given in [28].
More recently, [40] introduced a notion of ‘consistent Bayesian interpretation.’ This is a slightly different version of the ‘as-if’ approach, in which the agent’s internal states are interpreted as having a semantic meaning, rather than only its externally observable behaviour. It introduces an important element, which is that it makes sense to ask for an ‘as-if’ interpretation to be consistent over time. In particular, if we can treat a system as if it has a Bayesian prior at time $t$, then at time $t+1$, after having received some new sensory information, we should be able to treat the system as having a new set of beliefs, given by the appropriate Bayesian posterior. Ref. [40] shows that the conditions for such a ‘consistent Bayesian interpretation’ to be possible are relatively weak, at least in formal terms. When such interpretations exist they are not unique, which is an important aspect of the ‘as-if’ approach in general: there is rarely only one way in which a system can reasonably be treated as if it were an agent. This approach is extended to rational behaviour using value functions in [5].
In the current paper we are also concerned with consistency. We show that we obtain a form of consistency (Section 5) if certain assumptions are met. The form of consistency is not what we might naively expect, since the update rule is given by ‘value-laden filtering’ (Definition 5.0.1), rather than vanilla Bayesian filtering. We also observe that this ‘value-laden filtering’ consistency does not hold in general in bounded-rational contexts. In particular, in Section 5.1, we describe a case in which it doesn’t hold, due to constraints on memory. In Section 6.5 we also discuss another consequence of optimality under memory constraints, namely that the uniquely optimal policy may be non-deterministic, as previously observed in the absent-minded driver problem ([32]) and discussed more generally in [21].
2.1 Computational mechanics and unifilar machines
On a technical level our work is related to the tools of computational mechanics [36]. As in computational mechanics we are concerned with the externally observable behaviour of stochastic input-output processes, and the way in which we do this has something in common with the idea of an $\varepsilon$-transducer [3]. The formal objects known as $\varepsilon$-machines and $\varepsilon$-transducers are the minimal unifilar representations of stochastic processes (usually assumed to be stationary), and their states are called causal states. The concept of unifilarity plays an important role in our work (see Section 3.3.2). This concept originally comes from the computational mechanics literature, since it is used in one of the ways that one can define an $\varepsilon$-machine [38], or more generally an $\varepsilon$-transducer [3]. However, in our case, instead of using a minimal unifilar representation we are concerned with the equivalent of causal states as mathematical objects in their own right (our transducers). The connection between these ideas is made more precise in [39], where transducers are shown to form the states of a single unifilar machine. (The terminal one, in the sense of category theory.) We also note that the mapping from stochastic Moore machines to unifilar machines (our definition 3.3.7) has much in common with the ‘mixed state representation’ of a stochastic processes described in [8].
2.2 Representation Theorems For Expected Utility Theory
It is helpful to compare our approach to classic ‘representation theorems’ for expected utility theory from the decision-theoretic literature. A representation theorem (for expected utility) shows that, if a preference ordering $\prec$ over ‘prospects’ $O$ has certain formal properties, then it is equivalent to a ranking over prospects according to their expected value. As explained in [37] there are three well-known representation theorems for expected utility: due to von Neumann / Morgenstern (VNM) [29], Savage [35], and Bolker (used in Bolker-Jeffrey decision theory). In VNM the prospects (the objects of the preference order) are ‘lotteries’; in Savage they are ‘acts’; and in Bolker they are ‘propositions’.
In some loose sense, we provide, in this paper, a representation result for deterministic behavioural propensities: we observe that they can be represented as the uniquely optimal solutions to control problems. However, the domain in which we are working is very different from that of decision theory. We suppose that the object to be represented is the behaviour of a system; unlike VNM, Savage or Bolker/Jeffrey, we do not assume that information about an agent’s preferences are directly available to the theorist. This is for reasons similar to those described in [19]: behaviour (the agent’s ‘choices’) does not reveal an agent’s preferences unless we know the agent’s beliefs, and beliefs in turn cannot be inferred from behaviour without information about the agent’s evaluative judgements (which preferences are supposed to capture).
As discussed by [44], the classic ‘representations’ of preference orderings are non-unique, even given a fixed form of representation. Ours is no exception: there may be, and in general are, many different teleo-environments for which a given transducer is uniquely optimal. However, in the classical representation frameworks, the representations are unique up to a class of fairly simple transformations. This does not seem to be the case for our framework; there may be many teleo-environments that specify a given transducer, and these can be of very different character.
Another related family of classical results are the complete class theorems, for example [41, 6]. These have the general character that, under differing assumptions, a decision rule with a property called being non-dominated also has the property of being ‘Bayes-optimal’, meaning roughly that it can be thought of as doing loss minimisation with respect to some prior over the ‘states of nature.’ In our framework there is no obvious analog of the states of nature, so our results are of a different character; nevertheless it might be interesting to explore relationships in future work.
2.3 The free energy principle
An important contemporary area of research that touches on assigning mental states to systems is the free energy principle [16, 15], also known as Bayesian mechanics [33]. This is a broad body of theory that largely seeks applications in neuroscience. However, it also applies to much more general physical systems. Its proponents have talked about assigning mental states to non-biological physical systems since the earliest papers, e.g. [16]; this is made very explicit in [15]. This body of work tends to use a different formal approach from the current work, focused on stochastic differential equations rather than stochastic processes over discrete alphabets. Aside from technical differences, there are a few more foundational differences. The free energy principle is concerned with internal states rather than only externally observable behaviour. Perhaps more significantly, work on the free energy principle tends to seek a single, privileged belief state for a given system, which map states of the system to probability measures that relate to an ‘actual’ or typical environment in which the system is found. In the current paper we are focused on all possible assignments of belief states, including those in which the system might believe it inhabits a very different environment from the one it actually does; for this reason, we do not consider the ‘true’ environment at all in the current work.
2.4 Inverse Reinforcement Learning
Our work has some obvious similarities to an idea in machine learning known as inverse reinforcement learning. Inverse reinforcement learning is the problem of inferring the reward function of an agent, given its policy or observed behavior [2]. Generally, solutions to these problems assume that information about the environment is given, either explicitly in the form of a transition function, or implicitly in the form of trajectories. There are some partial exceptions to this, which attempt to infer the reward function with much more limited knowledge of the environment [7, 24], but none that attempt to infer the whole mental state of the agent based purely on its policy, which would be the true equivalent of our approach. This is not surprising, as inverse reinforcement learning is famously an already underspecified problem suffering from high computation costs, so the equivalent of our approach would likely be computationally infeasible. We don’t have such a constraint because we only aim to consider the abstract formal properties of such systems, rather than computing practical examples.
This concludes our review of previous work; in the next section we begin the main body of the paper.
3 Transducers
To mathematically describe the environments and policies we will be discussing in this paper we will use the basic notion of a transducer. Conceptually, as mentioned in the introduction, a transducer is something that takes a sequence of inputs and returns a sequence of outputs, which may depend stochastically on the inputs. This dependence must be causal, in the sense that a given output can only depend on inputs that were received at earlier times.
However, in order to reason formally about transducers we will instead think of them in a different way. We think of a transducer as something that can return an output stochastically. Having received its output we can then give it an input, at which point the transducer will transform into a new transducer, which will in general have a different probability distribution for its next output. In this way we can give the transducer a sequence of inputs and receive a sequence of outputs, and the causality condition will be satisfied automatically.
We will employ this notion of transducer in several different ways. We will use transducers to model the input-output behaviour of the systems we are interested in, but we will also use them to model those agents’ supposed beliefs about their environment. It will turn out that we can also use transducers to represent stochastic mixtures of other transducers - we will explain this concept below.
Our concept of transducer has something in common with the notion of $\varepsilon$-transducer defined in [3], in that both model a stochastic process that takes a sequence of inputs and generates a sequence of outputs, with each output being able to depend on inputs that precede it. However, the $\varepsilon$-transducer framework is concerned with stationary processes, which leads to a substantially different formalism. Our transducers also have much in common with controlled stochastic processes and other similar concepts from artificial intelligence and control theory. For example, the chronological semimeasures in [20] are an equivalent mathematical object and are also used to model the policies and environments of agents. Our main reason for coining a new term is our focus on the coinductive description, and in particular on the notion of evolving a transducer described in Section 3.1.
We will define transducers coinductively, as mentioned in Section 1.2. This allows us to use coinductive proofs, presented in the style proposed in [9]. We will also introduce some additional notation particularly useful for this style.
Now on to define our main object of interest. In the following, given a finite set $S$, we write $P(S)$ for the set of all probability distributions over $S$, that is, the set of functions $p\colon S\to[0,1]$ such that $\sum_{a\in A}p(a)=1$. Given a probability distribution $p\in P(S)$, we write $\mathrm{supp}(p)$ for the support of $p$, which is a subset of $S$.
Definition 3.0.1 (Transducer).
Let $O$ and $I$ be finite sets called the output space and input space respectively. We define the set of transducers from $I$ to $O$, written $I\mathrel{\triangleright}O$, coinductively using the single constructor taking $p\in P\left(O\right)$ and $t\colon I\times\mathrm{supp}\left(p\right)\to I\mathrel{\triangleright}O$ written as $\left(p,t\right)$.
Given a transducer $\pi\colon I\mathrel{\triangleright}O$ we will write $\mathbb{P}_{\pi}$ for the probability distribution in the first argument of the constructor, and $\mathbb{T}_{\pi}$ for the transition function in the second argument. The functions $\mathbb{P}$ and $\mathbb{T}$ are sometimes called destructors.
Briefly, this means that a transducer $\pi$ is an object that provides two things: first, it provides a way to stochastically generate a member of the output set, given by the distribution $\mathbb{P}_{\pi}$. Second, it provides a way to obtain a new transducer, given an input $i\in I$ and an output $o\in\mathrm{supp}(\mathbb{P}_{\pi})$. We can think of a transducer as specifying a conditional distribution over output sequences given input sequences. If we have given the transducer an input $i$ and we have observed that it produced the output $o$ in response, then we can calculate a new conditional distribution over output sequences given input sequences, starting from the next time step. The result is again a transducer, and in fact it is the one given by $\mathbb{T}_{\pi}(i,o)$. The reason the transition function takes an element of $\mathrm{supp}(\mathbb{P}_{\pi})$ rather than $O$ is that we can only calculate conditional distributions for those outputs that can occur with positive probability. We explore this connection more in section 3.2, where we also show that this operation can be seen as a form of Bayesian conditioning.
For readers familiar with the category-theoretic approach to coalgebra as in [22], a transducer is an element of the final coalgebra of the polynomial functor $\sum_{p\in P(O)}(I\times\mathrm{supp}(p)\to{-})$, or $\sum_{p\in P(O)}y^{I\times\mathrm{supp}(p)}$ in the notation of [30]. An explicit construction of this terminal coalgebra is given in [39]. However, we will generally not mention the category-theoretic approach, preferring instead to reason in the style proposed in [9], resembling classical inductive reasoning.
If we have a countable set of probability distributions, together with a probability weight assigned to each one, we can form a mixture distribution, which is itself a probability distribution. In the same way, if we have a sequence of transducers together with a sequence of probability weights assigned to them, we can form a mixture of the transducers, which is itself a transducer. One application for this is the case where we are dealing with an unknown transducer, drawn from a distribution with countable support.
Definition 3.0.2 (Mixture of transducers).
Given transducers $\pi_{k}\colon I\mathrel{\triangleright}O$ and numbers $\alpha_{k}\geq 0$, such that $\sum_{k=1}^{\infty}\alpha_{k}=1$, we define a (weighted) mixture $\sigma=\sum_{k=1}^{\infty}\alpha_{k}\pi_{k}$ as a transducer $(\mathbb{P}_{\sigma},\mathbb{T}_{\sigma})$ with
$\mathbb{P}_{\sigma}=\sum_{k=1}^{\infty}\alpha_{k}\mathbb{P}_{\pi_{k}}$ (1)
and
$\mathbb{T}_{\sigma}(i,o)=\sum_{k=1}^{\infty}\frac{\alpha_{k}\mathbb{P}_{\pi_{k}}\left(o\right)}{\sum_{l=1}^{\infty}\alpha_{l}\mathbb{P}_{\pi_{l}}\left(o\right)}\mathbb{T}_{\pi_{k}}\left(i,o\right),$ (2)
for $i\in I,o\in\mathrm{supp}(\mathbb{P}_{\sigma})$.
Note that this is a coinductive definition, since Equation 2 uses the notion of a mixture of transducers on its right-hand side.
When only a finite number of the coefficients are nonzero we will write these mixtures as ordinary sums, omitting the zero terms.
It is not a coincidence that Equation 2 resembles a Bayesian update. A mixture of transducers can be interpreted in Bayesian terms as a prior over these transducers (or technically, over the indices of a list of transducers) with the coefficients being the associated probabilities. If we feed the unknown transducer an input $i$ it will emit an output $o$ and evolve into a new transducer. When we receive the output $o$ we must use Bayes’ theorem to update our prior to a posterior, which is what Equation 2 expresses.
3.1 Evolutions
Intuitively a transducer $\pi\colon I\mathrel{\triangleright}O$ outputs an element $o\in O$ with probability $\mathbb{P}_{\pi}\left(o\right)$ and receives an input element $i\in I$, turning into the transducer $\mathbb{T}_{\pi}\left(i,o\right)$ in the process. This operation is the basis of all the behaviours we investigate in this paper, and one of the main reasons for choosing the definitions we are working with is how naturally it arises within this context.
In this section we extend this operation from one time step to many and introduce notation for it, as well as some related necessary notation. We start with the latter.
Definition 3.1.1 (Notation for strings).
Given a set $A$ we will consider the set of finite strings of elements of this set $A^{*}$. We will write $a,b,\dots$ for elements of $A$ and $\mathbf{a},\mathbf{b},\dots$ for elements of $A^{*}$.
We will denote concatenation of two strings by $\mathbf{a}\cdot\mathbf{b}$, overloading the operation for single elements as well, $\mathbf{a}\cdot b$. We also denote the empty string as $\varepsilon$.
We denote the element of a string at position $n$ by $\mathbf{a}_{n}$ (with the indexing starting from 0), and prefixes and suffixes as $\mathbf{a}_{<n}$ and $\mathbf{a}_{>n}$ respectively.
Given two strings of the same length $\mathbf{a}\in A^{n},\mathbf{b}\in B^{n}$ we write $\left(\mathbf{a}:\mathbf{b}\right)\in\left(A\times B\right)^{n}$ for the interleaved string $\left(\left(\mathbf{a}_{i},\mathbf{b}_{i}\right)\right)_{i=0}^{n}$. We will omit any additional parentheses when working with more than two strings, just as we do with tuples.
With this background we can define transducer evolutions.
Definition 3.1.2 (Evolution).
We define the operation of transducer evolution by finite trajectories using the infix operator ${}\bullet\colon I\mathrel{\triangleright}O\to\left(I\times O\right)^{*}\to I\mathrel{\triangleright}O\cup\left\{\star\right\}$ recursively with the base
$\pi\bullet\left(\varepsilon:\varepsilon\right)=\pi,$ (3)
and the step
$\pi\bullet\left(\mathbf{i}\cdot i:\mathbf{o}\cdot o\right)=\begin{cases}\mathbb{T}_{\pi\bullet\left(\mathbf{i}:\mathbf{o}\right)}\left(i,o\right)&\textrm{ if }\pi\bullet\left(\mathbf{i}:\mathbf{o}\right)\neq\star\textrm{ and}o\in\mathrm{supp}\left(\mathbb{P}_{\pi\bullet\left(\mathbf{i},\mathbf{o}\right)}\right),\\ \star&\textrm{ otherwise.}\end{cases}$ (4)
We will refer to an evolution returning a transducer (as opposed to $\star$) as ‘valid’.
3.2 Unrolling transducers
We present an alternative way of viewing transducers, that might improve intuitions and serves as a justification for some naming later. In particular, it clarifies the relationship between transducers and Bayes’ rule, which plays a fundamental role when transducers are formulated this way.
In this section, given a function $q\colon X\to P\left(Y\right)$ where $X$ and $Y$ are finite sets, we will write $q(y\parallel x)$ for the probability of the outcome $y$ according to the probability distribution $q(x)$, to stress the interpretation as conditional probability.
One can think of a transducer as something that takes in a sequence of inputs and stochastically generates a sequence of outputs, subject to the constraint that each element of the output sequence can only depend on the inputs that precede it in time. To do this we will show that the following definition is equivalent to definition 3.0.1.
Definition 3.2.1 (Unrolled transducers).
An unrolled transducer from $I$ to $O$ consists of a family of functions $p_{n}\colon I^{n}\to P\left(O^{n+1}\right)$, one for each $n\in\mathbb{N}$. These must have the property that
$\sum_{o\in O}p_{n+1}(\mathbf{o}\cdot o\parallel\mathbf{i}\cdot i)=p_{n}(\mathbf{o}\parallel\mathbf{i}),$ (5)
[…]
In other words, for every finite sequence of inputs, an unrolled transducer specifies a probability distribution over finite sequences of outputs. These output sequences are longer than the input sequences because we use the convention that a transducer gives its first output before it receives any input.
The condition in equation (5) serves two purposes: firstly it says that these conditional probability distributions of output sequences must all be consistent with one another, and secondly it says that the marginal probability distribution over outputs up to time $n+1$ can only depend on the inputs up to time $n$. This is the causality condition mentioned previously: outputs may depend on past inputs, but not on future ones.
Conceptually, an unrolled transducer specifies conditional probabilities of infinite output sequences given infinite input sequences, subject to a set of conditional independence relationships induced by the causality condition. However, we avoid the use of measure theory by talking about probability distributions over finite sequences of every length instead of measures over infinite sequences. This could be justified using a version of Kolmogorov’s extension theorem, which says that knowing a probability distribution over every finite sequence of symbols (such that they agree on all the marginals) is equivalent to knowing a probability measure over the space of all infinite sequences. We will never need to explicitly take such a step though, since by working coinductively we only ever need to consider finitely supported distributions over one symbol at a time.
[…]
Definition 3.2.2.
[…]
$u_{0}\left(p,t\right)(o)=p(o),$ (6)
and the step
$u_{n+1}\left(p,t\right)(o\cdot\mathbf{o}\parallel i\cdot\mathbf{i})=p(o)u_{n}(t(i,o))(\mathbf{o}\parallel\mathbf{i}).$ (7)
As we will often do, we gloss over the transition sometimes not being defined, since this is only the case when the probability would be zero anyway.
To see that this definition is correct, i.e. the resulting family of probabilities satisfies equation (5), we reason by induction. For the base we have
$\sum_{o^{\prime}\in O}u_{1}(p,t)(o\cdot o^{\prime}\parallel i)=\sum_{o^{\prime}\in O}p(o)\mathbb{P}_{t(i,o)}(o^{\prime}),$ (8)
where the right hand side sums to $p(o)=u_{0}(p,t)(o)$ as desired. Similarly for the inductive step
$\sum_{o^{\prime}\in O}u_{n+1}(p,t)(o\cdot\mathbf{o}\cdot o^{\prime}\parallel i\cdot\mathbf{i}\cdot i^{\prime})=\sum_{o^{\prime}\in O}p(o)u_{n}(t(i,o))(\mathbf{o}\cdot o^{\prime}\parallel\mathbf{i}\cdot i^{\prime}),$ (9)
whence we can apply the inductive hypothesis
$p(o)u_{n-1}(t(i,o))(\mathbf{o}\parallel\mathbf{i}),$ (10)
which is exactly the definition of
$u_{n}(p,t)(o\cdot\mathbf{o}\parallel i\cdot\mathbf{i}),$ (11)
as desired.
It remains to define an inverse and show that it is such. To do this it’s useful to first define a transition function for unrolled transducers.
Definition 3.2.3.
The transition function for unrolled transducers
[…]
is defined as
$s(p,i,o)_{n}(\mathbf{o}\parallel\mathbf{i})=\frac{p_{n+1}(o\cdot\mathbf{o}\parallel i\cdot\mathbf{i})}{p_{0}(o)}.$ (12)
To see that the transition function for unrolled transducers actually returns an unrolled transducer, note that it just represents conditioning all the functions in the family on the first output being $o$ and the first input being $i$, which preserves condition (5).
Note that the transition function for unrolled transducers is nothing but Bayesian conditioning on an input-output pair. This fact will allow us to treat evolutions of transducers as Bayesian updates in the following. We just need to finish proving the equivalence.
Lemma 3.2.4.
The function $u$ has an inverse defined coinductively as
$r(p)=\left(p,(i,o)\mapsto r(s(p,i,o)\right),$ (13)
Proof.
We will first prove that the first composition $u(r(p))$ is the identity by induction, starting with
$u_{0}(r(p))=p_{0},$ (14)
and the step
$u_{n+1}(r(p))=u_{n+1}\left(\left(p_{0},(i,o)\mapsto r(s(p,i,o))\right)\right),$ (15)
which on specific arguments $o\cdot\mathbf{o}\in O^{n+2}$, $i\cdot\mathbf{i}\in I^{n+1}$ is
$p_{0}(o)u_{n}(r(s(p,i,o)))(\mathbf{o}\parallel\mathbf{i}),$ (16)
letting us apply the inductive hypothesis
$p_{0}(o)\frac{p_{n+1}(o\cdot\mathbf{o}\parallel i\cdot\mathbf{i})}{p_{0}(o)}=p_{n+1}(o\cdot\mathbf{o}\parallel i\cdot\mathbf{i}),$ (17)
as required.
Now let us examine the other composition
$\left(u_{0}(p,t),(i,o)\mapsto r(s(u(p,t),i,o)))\right).$ (18)
Using the definition of $u_{0}$ and noting that the multiplication and division in definitions of $s$ and $u$ cancel out we get
$\left(p,(i,o)\mapsto r(u(t(i,o)))\right),$ (19)
so it remains to use the coinductive hypothesis to finish the proof. ∎
Since this is the first time we are using this style of reasoning, let us explain it in a bit more detail. We rely on a ‘coinductive hypothesis’ that the relevant equality holds for any ‘smaller’ or ‘deconstructed’ object. In the case of the above proof we acquire such an object in the form of $t(i,o)$ – it is ‘smaller’, intuitively, due to us discarding some information, in the form of $p$, from the full object. This principle can be applied not only to equality but also to properties known as ‘coinductively defined’ properties; we make use of this in later sections. A reader interested in a more complete formal explanation of the above can find it in [9, Example 4.30].
3.3 Constrained transducers
In order to have a hope of being applied to practical situations it is important to take into account the computational constrains of real world systems, as argued in [42, Section 4.1]. To this end we will sometimes consider constrained classes of transducers, which in principle are arbitrary subsets of all transducers $T\subseteq I\mathrel{\triangleright}O$, but in practice will be somewhat less arbitrary. In this section we will describe some properties such classes can have and define some particularly useful ones.
3.3.1 Exchange property
In this section we introduce the exchange property, which we will later use as an assumption in our main theorem, and illustrate this property with some examples.
Definition 3.3.1 (Exchange property).
We say that a set of constrained transducers $T\subseteq I\mathrel{\triangleright}O$ has the exchange property if, given $\pi,\pi^{\prime}\in T$ and an input-output pair $\left(i,o\right)\in I\times O$ valid for evolving $\pi$, the transducer defined as
$\left(\mathbb{P}_{\pi},\left(j,q\right)\mapsto\begin{cases}\pi^{\prime}&\textrm{ when }\left(j,q\right)=\left(i,o\right),\\ \pi\bullet\left(j:q\right)&\textrm{ otherwise,}\end{cases}\right)$ (20)
also belongs to $T$.
In other words, the operation of replacing the future transducer after a specific input-output pair by an arbitrary one from the same class is well defined within the class. In particular, this implies that if a class can be defined purely coinductively, i.e. by constructing its elements so that they can evolve into an arbitrary member of this class, then it does have the exchange property. This includes the class of all transducers.
The exchange property is equivalent to a property that replaces the future transducer after evolving by an arbitrary trajectory, a fact that is straightforward to prove inductively. We won’t use this fact explicitly in any proofs, but it might help build intuition.
Example 3.3.2 (i.i.d. transducers).
An independent and identically distributed transducer or i.i.d. transducer is one that can be defined as
$\pi=\left(p,\left(i,o\right)\mapsto\pi\right),$ (21)
for some $p\in P\left(O\right)$. Note that while a single i.i.d. transducer is defined coinductively, the whole class cannot be.
i.i.d. transducers do not have the exchange property if the support of $p$ has more than one element, since if we replaced the future transducer by another i.i.d. one the resulting transducer would no longer be i.i.d.
Example 3.3.3 (Deterministic transducers).
A deterministic transducer $\pi\colon D\subset I\mathrel{\triangleright}O$ is one for which the associated probability distribution is a one-point distribution and the transition function only takes values in further deterministic transducers $D$.
This is a purely coinductive definition, so deterministic transducers have the exchange property.
We will also sometimes refer to intersections of this class with other classes by adding the adjective ‘deterministic’ to the name of the given class.
More generally any constraint on probability distributions independent of any transducer structure (e.g. minimal probability assigned to all possibilities, uniform over a subset, etc.) gives rise to a class of transducers with the exchange property. However, this is not a necessary condition, as the following class demonstrates.
Example 3.3.4 (Smearing transducers).
A smearing transducer of order $k\in\mathbb{N}$ has an associated probability with a support with at most $k$ elements, and its transition function takes values in smearing transducers of order $k+1$.
Smearing transducers have the exchange property, because the ones of order $k$ are contained among the ones of any higher order.
Example 3.3.5 (One-flip transducers).
A one-flip transducer is one for which the associated probability distribution either is a one-point distribution and the transition function takes values within one-flip transducers, or an arbitrary distribution and the transition function takes values in deterministic transducers. Thus these transducers can have at most one associated random variable not be deterministic (they have ‘one coin flip’, hence the name).
One-flip transducers don’t have the exchange property, because once they use their coin flip they cannot have a transition lead to a one-flip transducer that has not yet used the coinflip. They will also be an important example for another reason later.
3.3.2 Unifilar machines
In this and the following sections we introduce several formalisms for modelling state-based dynamics, which can easily be translated into transducers. We start with unifilar machines, as they are used as an intermediate step in later sections.
Definition 3.3.6 (Unifilar machines).
Let $X$ be an arbitrary set called the state set. We will call $\left(\varphi,\theta\right)\colon X\to\sum_{\varphi\colon P\left(O\right)}I\times\mathrm{supp}\left(\varphi\right)\to X$ a unifilar machine. Given a state $x\colon X$ we will write $\left(\varphi_{x},\theta_{x}\right)=\left(\varphi,\theta\right)\left(x\right)$ and call it a pointed unifilar machine. The elements of the pair $\left(\varphi,\theta\right)$ will be called its output function, and transition function respectively. We denote the whole set of unifilar machines over $X$ as $U_{X}$.
Unifilar machines take their name from the similar concept of unifilarity defined in the context of $\varepsilon$-transducers in [3].
A reader familiar with more explicit treatments of coinduction may notice that a unifilar machine is a coalgebra of the functor $\sum_{\varphi\colon P\left(O\right)}I\times\mathrm{supp}\left(\varphi\right)\to{-}$, while, as already mentioned, the set of all transducers, $I\mathrel{\triangleright}O$, is the carrier of the final coalgebra of this functor. Our coinductive definition is equivalent to specifying the final coalgebra of this functor. This will not make any significant difference below, as we explicitly define a map into transducers anyway.
Definition 3.3.7.
Define a mapping $s\colon X\times U_{X}\to I\mathrel{\triangleright}O$ from pointed unifilar machines to transducers corecursively as
$s\left(x,\left(\varphi,\theta\right)\right)=\left(\varphi_{x},\left(i,o\right)\mapsto s\left(\theta_{x}\left(i,o\right),\left(\varphi,\theta\right)\right)\right).$ (22)
3.3.3 Unifilar finite state machines
We will sometimes want to consider certain subsets of transducers, corresponding to transducers that can be constructed in a particular way. In this section we will see our first example. We consider a special case of unifilar machines, which represents a formalism for limited memory (but unlimited time) computation. We then discuss the transducers that can be constructed from these machines.
Definition 3.3.8 (Unifilar finite state machines).
Given a nonempty finite state set of size $n$ $X$, we will call $\left(\varphi,\theta\right)\in U_{X}$ a unifilar finite state (UFS) machine.
We interpret this state set as all the possible states of memory, and output and transition functions as code that maps a memory state to a probability distribution over outputs and a memory state plus an input and output into a new memory state respectively. In this abstraction the computational process has access to a randomness source.
Unifilar finite state machines are roughly equivalent to deterministic finite state (Moore) machines, although some of the considered sets are infinite in contrast to the classical definition.
We can now map these unifilar transducers into transducers using $s$ defined in the previous section. As one might expect the state set will become irrelevant through this mapping, so we will encounter the problem that there are multiple distinct UFS machines that give rise to the same transducers (e.g. by composing the machine with a bijection to another set), so we will have to keep that in mind when discussing examples.
Definition 3.3.9 (Unifilar finite state transducers).
We call the image $\mathrm{UFS}_{n}=s\left(X\times U_{X}\right)$ unifilar finite state (UFS) transducers.
It is easy to see that for all $n\in\mathbb{Z}_{+}$ $\mathrm{UFS}_{n}\subseteq\mathrm{UFS}_{n+1}$. It suffices to notice that the subset of UFS machines that identifies two elements of the memory set (i.e the machine behaves identically regardless of which of these elements is passed to it) gives rise to the UFS transducers with one less possible memory state.
It is similarly easy to see that for $n=1$ UFS transducers are exactly i.i.d. transducers, so they don’t have the exchange property if the output set has more than one element. More generally no UFS transducers have the exchange property if that’s the case.
3.3.4 Stochastic Moore machines
In this section we introduce yet another notion of state-based dynamics. We won’t be using any object constructed here in the rest of the paper, but they serve as important motivation and context for interpreting the main result.
This is the only section explicitly using measure theory. For the most part, a reader unfamiliar with measure theory may ignore the measure theoretic details. In previous sections we used $P\left(A\right)$ to mean the set of probability distributions over a finite set $A$. In this section we extend this notation, so that $P\left(Y\right)$ means the space of probability measures over a measurable space $Y$, where $P\left(Y\right)$ itself is to be viewed as a measurable space, equipped with the smallest $\sigma$-algebra such that for every event in the $\sigma$-algebra of $S$, the map $P\left(Y\right)\to[0,1]$ given by $\mu\mapsto\mu(S)$ is measurable. In category theory terms this means that in this section $P$ stands for the functor underlying the Giry monad [18, 17]. When $X$ is a finite set, $P\left(X\right)$ still amounts to the set of probability distributions over $X$. We will continue to assume $I$ and $O$ are finite sets.
Recall that for measurable spaces $X$ and $Y$, a Markov kernel from $X$ to $Y$ is a measurable function $X\to P\left(Y\right)$. Given a Markov kernel $q\colon X\to P\left(Y\right)$ in which $Y$ is finite, we will write $q(y\parallel x)$ for the probability of the outcome $y$ according to the probability distribution $q(x)$, as in section 3.2.
We first define a basic notion of a machine with an internal state:
Definition 3.3.10 (Stochastic Moore machines).
A stochastic Moore machine consists of a measurable space $Y$ called the state space, together with a Markov kernel $\varphi\colon Y\to P\left(O\right)$ called the output kernel, a Markov kernel $\theta\colon Y\times I\to P\left(Y\right)$ called the transition kernel and a probability measure $\psi\in P\left(Y\right)$ called the initial distribution. We call the triple $\left(\psi,\varphi,\theta\right)$ a stochastic Moore machine over $Y$. We denote the set of stochastic Moore machines over $Y$ as $M_{Y}$.
The idea is that a stochastic Moore machine has an internal state, which is an element of $Y$. It produces an output stochastically according to its output kernel. Then, after being given an input, it transitions stochastically to a new state, according to its transition function. In general the state of a stochastic Moore machine might not be known. (And even if it is initially known, it will generally be in an unknown state after a transition, assuming the observer can’t see inside the machine.) We define a stochastic Moore machine to be equipped with an initial probability measure over its state space, to be thought of as a prior.
In contrast to unifilar machines, which can only introduce nondeterminism via the output kernel, stochastic Moore machines can have nondeterministic transitions that cannot depend on the output. Despite this difference, given a stochastic Moore machine there is a way to define a unifilar machine with the same behaviour:
Definition 3.3.11.
Given a state space $Y$, we can map stochastic Moore machines over $Y$ into pointed unifilar machines over $P\left(Y\right)$. We denote this map
$z\colon M_{Y}\to\left(P\left(Y\right)\times U_{P\left(Y\right)}\right).$
It is defined on the components as
$z\left(\psi\right)=\psi,$ (23)
on the state mixture, with the following operation on the output kernels
$z\left(\varphi\right)\left(\psi\right)=\int_{Y}\varphi\left(y\right)\psi\left(y\right)\mathrm{d}y,$ (24)
and the somewhat more complex operation on transition kernels
$z\left(\theta\right)\left(\psi,i,o\right)=\frac{1}{z\left(\varphi\right)\left(o\parallel\psi\right)}\int_{Y}\theta\left(y,i\right)\varphi\left(o\parallel y\right)\psi\left(y\right)\mathrm{d}y.$ (25)
A remark is in order on the meaning of this map. On the one hand, intuitively, the pointed unifilar machine $z(\psi,\varphi,\theta)$ ‘behaves the same’ as the stochastic Moore machine $(\psi,\varphi,\theta)$, but on the other hand it is a different type of machine and has a different state space. The states of the unifilar machine consist of probability distributions over $Y$ instead of elements of $Y$. One way to think of this is that these distributions represent an agent’s state of knowledge about the state of the corresponding Moore machine. Initially this state of knowledge is given by the stochastic Moore machine’s initial distribution. Then equation (24) can be seen as a prediction of the Moore machine’s next output, and (25) can be seen as a Bayesian update, producing a posterior over the Moore machine’s states, conditioned on the given input and the observed output.
Remark 3.3.12.
This updating of a probability distribution over $Y$ may be seen as an instance of Bayesian filtering with an additional input [39]. This perspective will be further justified in Section 4.1.1.
Now we can use the composition $s\circ z$ to map stochastic Moore machines into transducers. As one would expect, this mapping respects mixtures in the first argument, we will show that in two steps.
Lemma 3.3.13.
Suppose we are given a machine $\left(\varphi,\theta\right)\colon U_{P\left(Y\right)}$ for which the output function respects mixtures and the transition function satisifes
$\theta_{\sum_{k=1}^{\infty}\alpha_{k}\psi_{k}}\left(i,o\right)=\sum_{k=1}^{\infty}\frac{\alpha_{k}\varphi_{\psi_{k}}\left(o\right)}{\sum_{l=1}^{\infty}\alpha_{l}\varphi_{\psi_{l}}\left(o\right)}\theta_{\psi_{k}}\left(i,o\right),$ (26)
for any numbers $\alpha_{k}\geq 0$ such that $\sum_{k=0}^{\infty}\alpha_{k}=1$. Then for any such numbers $s$ respects mixtures of (distributions of) states for an unifilar machine with these output and transition functions, that is
$s\left(\sum_{k=1}^{\infty}\alpha_{k}\psi_{k},\left(\varphi,\theta\right)\right)=\sum_{k=1}^{\infty}\alpha_{k}s\left(\psi_{k},\left(\varphi,\theta\right)\right).$ (27)
Proof.
Starting with the left hand side and expanding the definition of $s$
$s\left(\sum_{k=1}^{\infty}\alpha_{k}\psi_{k},\left(\varphi,\theta\right)\right)=\left(\varphi_{\sum_{k=1}^{\infty}\alpha_{k}\psi_{k}},\left(i,o\right)\mapsto s\left(\theta_{\sum_{k=1}^{\infty}\alpha_{k}\psi_{k}}\left(i,o\right),\left(\varphi,\theta\right)\right)\right).$ (28)
Applying the fact that $\varphi$ respects mixtures and the property from equation (26) we get
$\left(\sum_{k=1}^{\infty}\alpha_{k}\varphi_{\psi_{k}},\left(i,o\right)\mapsto s\left(\sum_{k=1}^{\infty}\frac{\alpha_{k}\varphi_{\psi_{k}}\left(o\right)}{\sum_{l=1}^{\infty}\alpha_{l}\varphi_{\psi_{l}}\left(o\right)}\theta_{\psi_{k}}\left(i,o\right),\left(\varphi,\theta\right)\right)\right).$ (29)
Using the coinductive hypothesis we can extract the mixture from the transition
$\left(\sum_{k=1}^{\infty}\alpha_{k}\varphi_{\psi_{k}},\left(i,o\right)\mapsto\sum_{k=1}^{\infty}\frac{\alpha_{k}\varphi_{\psi_{k}}\left(o\right)}{\sum_{l=1}^{\infty}\alpha_{l}\varphi_{\psi_{l}}\left(o\right)}s\left(\theta_{\psi_{k}}\left(i,o\right),\left(\varphi,\theta\right)\right)\right).$ (30)
This is exactly the definition of a mixture of transducers, giving us
$\sum_{k=1}^{\infty}\alpha_{k}\left(\varphi_{\psi_{k}},\left(i,o\right)\mapsto s\left(\theta_{\psi_{k}}\left(i,o\right),\left(\varphi,\theta\right)\right)\right).$ (31)
Collapsing the definition of $s$, gets us the desired
$\sum_{k=1}^{\infty}\alpha_{k}s\left(\psi_{k},\left(\varphi,\theta\right)\right).$ (32)
Lemma 3.3.14.
Given state distributions $\psi_{k}\colon P\left(X\right)$ and numbers $\alpha_{k}\geq 0$, such that $\sum_{k=0}^{\infty}\alpha_{k}=1$ we have
$s\circ z\left(\sum_{k=1}^{\infty}\alpha_{k}\psi_{k},\varphi,\theta\right)=\sum_{k=1}^{\infty}\alpha_{k}s\circ z\left(\psi_{k},\varphi,\theta\right).$ (33)
Proof.
We will prove that $z\left(\psi,\varphi,\theta\right)$ satisfies the conditions of Lemma 3.3.13 and, noting that it is the identity on the first component, an application of that lemma finishes the proof.
Let us start with the output function respecting mixtures
$z\left(\varphi\right)\left(\sum_{k=1}^{\infty}\alpha_{k}\psi_{k}\right)=\int_{Y}\varphi\left(y\right)\sum_{k=1}^{\infty}\alpha_{k}\psi_{k}\left(y\right)\mathrm{d}y\\ =\sum_{k=1}^{\infty}\alpha_{k}\int_{Y}\varphi\left(y\right)\psi_{k}\left(y\right)\mathrm{d}y=\sum_{k=1}^{\infty}\alpha_{k}z\left(\varphi\right)\left(\psi_{k}\right).$ (34)
Having this we can compute
$z\left(\theta\right)\left(\sum_{k=1}^{\infty}\alpha_{k}\psi_{k},i,o\right)=\frac{1}{\sum_{l=1}^{\infty}\alpha_{l}z\left(\varphi\right)\left(o\parallel\psi_{l}\right)}\int_{Y}\theta\left(y,i\right)\varphi\left(o\parallel y\right)\sum_{k=1}^{\infty}\alpha_{k}\psi_{k}\left(y\right)\mathrm{d}y$ (35)
writing $1=\frac{z\left(\varphi\right)\left(o\parallel\psi_{k}\right)}{z\left(\varphi\right)\left(o\parallel\psi_{k}\right)}$ and performing some simple operations on sums and integrals we get
$\sum_{k=0}^{\infty}\frac{\alpha_{k}z\left(\varphi\right)\left(o\parallel\psi_{k}\right)}{\sum_{l=1}^{\infty}\alpha_{l}z\left(\varphi\right)\left(o\parallel\psi_{l}\right)}\frac{1}{z\left(\varphi\right)\left(o\parallel\psi_{k}\right)}\int_{Y}\theta\left(y,i\right)\varphi\left(o\parallel y\right)\psi_{k}\left(y\right)\mathrm{d}y\\ =\sum_{k=0}^{\infty}\frac{\alpha_{k}z\left(\varphi\right)\left(o\parallel\psi_{k}\right)}{\sum_{l=1}^{\infty}\alpha_{l}z\left(\varphi\right)\left(o\parallel\psi_{l}\right)}z\left(\theta\right)\left(\psi_{k},i,o\right).$ (36)
It remains to apply Lemma 3.3.13 and we get the desired result. ∎
We perform the two steps separately to stress that only the first proof uses coinduction. The second could not have used it directly, as $z$ is not coinductively defined.
We conjecture that there is a $\sigma$-algebra on $I\mathrel{\triangleright}O$ that, among other things, would allow one to define maps in the other direction, from transducers (or subsets of transducers) to stochastic Moore machines. This in turn would allow some of the machinery of [39] to be applied in our context. The existence of such a $\sigma$-algebra is not obvious, as it requires several functions, both from and into transducers to be measurable. A suitable $\sigma$-algebra on the set of transducers would also allow reasoning about uncountable mixtures of transducers.
4 Teleo-Environments
We finally define the main objects we will be investigating.
Definition 4.0.1 (Teleo-environment).
Let $S$ and $A$ be nonempty finite sets called the state space and action space respectively. Furthermore let $G=\left\{\bot,\top\right\}$, called the telos channel, with the contents of the set called nothing and success respectively. We then call $\varepsilon\colon A\mathrel{\triangleright}\left(S\times G\right)$ a teleo-environment and $\pi\colon S\mathrel{\triangleright}A$ a policy.
One should think of ‘success’ and ‘nothing’ as follows: an agent’s goal in interacting with a teleo-environment is to maximise its probability of achieving success at least once. So $\bot$ does not represent failure, but only that success has not been achieved on this particular time step. The agent does not receive the success signal as an input, so in general the agent will not know whether success has been achieved yet or not; this fact will turn out to be important.
We remark that instead of maximising the probability of success, one could define a similar set-up where an agent tries to minimise a probability of failure; we would expect similar results to hold in that case. One could also consider an agent maximising a suitably bounded reward function, perhaps with exponential discounting, which is the more familiar set-up associated with POMDP tasks. However, this has an important difference, in that success (or failure) probabilities combine multiplicatively rather than additively on successive time steps, and for this reason we would expect the resulting formalism to be somewhat different. We will not consider either of these cases further.
There is a natural way to connect a policy and an environment together, getting a transducer that does not require inputs, which we will call a coupled system.
4.1 Coupling
Definition 4.1.1 (Coupling).
We define the coupling function
$c\colon\left(S\mathrel{\triangleright}A\right)\times\left(A\mathrel{\triangleright}\left(S\times G\right)\right)\to\{\star\}\mathrel{\triangleright}S\times G\times A$ (37)
by setting the probability distribution to the product distribution
$\mathbb{P}_{c\left(\pi,\varepsilon\right)}\left(s,t,a\right)=\mathbb{P}_{\pi}\left(a\right)\mathbb{P}_{\varepsilon}\left(s,t\right),$ (38)
and the transition function to a corecursive application of coupling to the transitions of the arguments
$\mathbb{T}_{c\left(\pi,\varepsilon\right)}\left(\star,\left(s,t,a\right)\right)=c\left(\mathbb{T}_{\pi}\left(s,a\right),\mathbb{T}_{\varepsilon}\left(a,\left(s,t\right)\right)\right).$ (39)
The $\star$ here is a dummy, and only possible, input to conform with the formal definition of a transducer.
The coupling operation will be crucial for our investigations, even though we will rarely refer to it explicitly. We will prove one important property it has though.
Lemma 4.1.2 (Coupling respects mixtures).
Let $\pi_{k}\colon S\mathrel{\triangleright}A$ be policies, $\varepsilon_{k}\colon A\mathrel{\triangleright}\left(S\times G\right)$ be environments, and $\alpha_{k},\beta_{k}\geq 0$. The coupling function respects mixtures, that is
$c\left(\sum_{k=1}^{\infty}\alpha_{k}\pi_{k},\sum_{k=1}^{\infty}\beta_{k}\varepsilon_{k}\right)=\sum_{k=1}^{\infty}\sum_{l=1}^{\infty}\alpha_{k}\beta_{l}c\left(\pi_{k},\varepsilon_{l}\right).$ (40)
Proof.
Let us start with the probability
$\mathbb{P}_{c\left(\sum_{k=1}^{\infty}\alpha_{k}\pi_{k},\sum_{k=1}^{\infty}\beta_{k}\varepsilon_{k}\right)}\left(s,t,a\right)=\mathbb{P}_{\sum_{k=1}^{\infty}\alpha_{k}\pi_{k}}\left(a\right)\mathbb{P}_{\sum_{k=1}^{\infty}\beta_{k}\varepsilon_{k}}\left(s,t\right).$ (41)
Applying the definition of a mixture this becomes
$\sum_{k=1}^{\infty}\alpha_{k}\mathbb{P}_{\pi_{k}}\left(a\right)\sum_{k=1}^{\infty}\beta_{k}\mathbb{P}_{\varepsilon_{k}}\left(s,t\right)=\sum_{k=1}^{\infty}\sum_{l=1}^{\infty}\alpha_{k}\beta_{l}\mathbb{P}_{\pi_{k}}\left(a\right)\mathbb{P}_{\varepsilon_{l}}\left(s,t\right),$ (42)
which after rewriting using the definition of coupling gives us exactly the definition of a probability of a mixture of couplings:
$\sum_{k=1}^{\infty}\sum_{l=1}^{\infty}\alpha_{k}\beta_{l}\mathbb{P}_{c\left(\pi_{k},\varepsilon_{l}\right)}\left(s,t,a\right).$ (43)
Now the transition function is given by
$\mathbb{T}_{c\left(\sum_{k=1}^{\infty}\alpha_{k}\pi_{k},\sum_{k=1}^{\infty}\beta_{k}\varepsilon_{k}\right)}\left(\star,\left(s,t,a\right)\right)=c\left(\mathbb{T}_{\sum_{k=1}^{\infty}\alpha_{k}\pi_{k}}\left(s,a\right),\mathbb{T}_{\sum_{k=1}^{\infty}\beta_{k}\varepsilon_{k}}\left(a,\left(s,t\right)\right)\right).$ (44)
Applying the definition of a mixture it becomes
$c\left(\sum_{k=1}^{\infty}\frac{\alpha_{k}\mathbb{P}_{\pi_{k}}\left(a\right)}{\sum_{l=1}^{\infty}\alpha_{l}\mathbb{P}_{\pi_{l}}\left(a\right)}\mathbb{T}_{\pi_{k}}\left(s,a\right),\sum_{k=1}^{\infty}\frac{\beta_{k}\mathbb{P}_{\varepsilon_{k}}\left(s,t\right)}{\sum_{l=1}^{\infty}\beta_{l}\mathbb{P}_{\varepsilon_{l}}\left(s,t\right)}\mathbb{T}_{\varepsilon_{k}}\left(a,s,t\right)\right).$ (45)
Using the coinductive hypothesis and performing simple operations on sums it becomes
$\sum_{k=1}^{\infty}\sum_{l=1}^{\infty}\frac{\alpha_{k}\beta_{l}\mathbb{P}_{\pi_{k}}\left(a\right)\mathbb{P}_{\varepsilon_{l}}\left(s,t\right)}{\sum_{k^{\prime}=1}^{\infty}\sum_{l^{\prime}=1}^{\infty}\alpha_{k^{\prime}}\beta_{l^{\prime}}\mathbb{P}_{\pi_{k^{\prime}}}\left(a\right)\mathbb{P}_{\varepsilon_{l^{\prime}}}\left(s,t\right)}c\left(\mathbb{T}_{\pi_{k}}\left(s,a\right),\mathbb{T}_{\varepsilon_{l}}\left(a,s,t\right)\right)$ (46)
and after rewriting using the definition of coupling it gives us exactly the definition of a transition of a mixture of couplings:
$\sum_{k=1}^{\infty}\sum_{l=1}^{\infty}\frac{\alpha_{k}\beta_{l}\mathbb{P}_{c\left(\pi_{k},\varepsilon_{l}\right)}\left(s,t,a\right)}{\sum_{k^{\prime}=1}^{\infty}\sum_{l^{\prime}=1}^{\infty}\alpha_{k^{\prime}}\beta_{l^{\prime}}\mathbb{P}_{c\left(\pi_{k^{\prime}},\varepsilon_{l^{\prime}}\right)}\left(s,t,a\right)}\mathbb{T}_{c\left(\pi_{k},\varepsilon_{l}\right)}\left(\star,\left(s,t,a\right)\right).$ (47)
4.1.1 Coupled evolutions as filtering
The term ‘Bayesian filtering’ is used to describe a technique in which Bayesian inference is performed sequentially, to track a variable which changes over time. In terms of random variables: upon receiving data $D_{n}$ at time step $n$, with a prior $P(X_{n}\mid D_{<n})$, filtering produces a posterior $P(X_{n+1}\mid D_{<n+1})$. Note that the posterior is over a different variable $X_{n+1}$ than the prior, which is over $X_{n}$. In this section we will explain why the evolution operator on coupled transducers can be understood in terms of filtering.
Since a coupling has trivial input, we can treat its unrolling as a probability space over output sequences, allowing us to refer to random variables within that space. We will write $P_{c\left(\pi,\varepsilon\right)}(S_{0},\cdots S_{n},T_{0},\cdots,T_{n},A_{0},\cdots,A_{n})$ for the probability space induced by the distribution $u_{n}\left(c\left(\pi,\varepsilon\right)\right)$. Then the evolution on couplings can be understood in terms of their unrollings as follows. For any $n,m\in\mathbb{N}$, any sequences of state data $\mathbf{s}\in S^{n},\mathbf{s^{\prime}}\in S^{m}$, any sequences of telos data $\mathbf{t}\in G^{n},\mathbf{t^{\prime}}\in G^{m}$, any sequences of action data $\mathbf{a}\in A^{n},\mathbf{a^{\prime}}\in A^{m}$ we have:
$P_{c\left(\pi\bullet\left(\mathbf{s}:\mathbf{a}\right),\varepsilon\bullet\left(\mathbf{a}:\left(\mathbf{s}:\mathbf{t}\right)\right)\right)}(S_{1\cdots m}=\mathbf{s^{\prime}},T_{1\cdots m}=\mathbf{t^{\prime}},A_{1\cdots m}=\mathbf{a^{\prime}})=\\ P_{c\left(\pi,\varepsilon\right)}(S_{n+1\cdots n+m}=\mathbf{s^{\prime}},T_{n+1\cdots n+m}=\mathbf{t^{\prime}},A_{n+1\cdots n+m}=\mathbf{a^{\prime}}\\ \mid S_{1\cdots n}=\mathbf{s},T_{1\cdots n}=\mathbf{t},A_{1\cdots n}=\mathbf{a})$ (48)
In other words, the first $m$ steps of the unrolling of the coupling of evolutions
$c\left(\pi\bullet\left(\mathbf{s}:\mathbf{a}\right),\varepsilon\bullet\left(\mathbf{a}:\left(\mathbf{s}:\mathbf{t}\right)\right)\right)$ (49)
behave like the steps $n+1\cdots n+m$ of the unrolling of the coupling $c\left(\pi,\varepsilon\right)$, conditioned on the first $n$ outputs being equal to the evolution trajectory. In this sense, the evolution operator can be seen as an analogue of Bayesian filtering.
Recall that Remark 3.3.12 pointed out the similarity between Bayesian filtering and the operation on Moore machines which corresponded to transition. Since evolution is just repeated transition we now see that the comparison was not coincidental – in the case of coupled systems this operation maps into filtering for unrollings.
4.2 Success
We now proceed to defining the probability of success. Note that we only care about encountering success at least once, multiple occurences are irrelevant.
We first define some helper objects and functions which will be crucial in proofs.
Definition 4.2.1 (Bounded sum sequences).
We define the family bounded sum sequences $\mathrm{BSS}\left(r_{0}\right)$ over the interval $\left[0,1\right]$ coinductively using the single constructor taking $r\in\left[0,r_{0}\right]$ and $t\colon\mathrm{BSS}\left(r_{0}-r\right)$ written as $\left(r,t\right)$.
The idea is that $\mathrm{BSS}\left(r_{0}\right)$ represents the set of sequences of numbers in $[0,1]$ that sum to at most $r_{0}$. Consequently when the first element is removed the remainder of the sequence sums to at most $r_{0}-r$, i.e. it is an element of $\mathrm{BSS}\left(r_{0}-r\right)$.
There are natural multiplication and summing operations on bounded sum sequences.
Definition 4.2.2.
Given numbers $r_{0},s\in\left[0,1\right]$ we define multiplication
$s\cdot\colon\mathrm{BSS}\left(r_{0}\right)\to\mathrm{BSS}\left(sr_{0}\right)$ (50)
corecursively as
$s\left(r,t\right)=\left(sr,st\right).$ (51)
Given numbers $s,s_{k}\in\left[0,1\right]$ such that $\sum_{i=0}^{\infty}s_{k}=s$ and a sequence of bounded sum sequences $\left(r_{k},t_{k}\right)\colon\mathrm{BSS}\left(s_{k}\right)$ we define summing corecursively as
$\sum_{k=0}^{\infty}\left(r_{k},t_{k}\right)=\left(\sum_{k=0}^{\infty}r_{k},\sum_{k=0}^{\infty}t_{k}\right),$ (52)
which belongs to $\mathrm{BSS}\left(s\right)$.
We can use bounded sum sequences to represent the probability that a specific coupled system outputs its first success exactly at a given step.
Definition 4.2.3 (Success sequence).
We define the success sequence $\sigma\colon\{\star\}\mathrel{\triangleright}S\times G\times A\to\mathrm{BSS}\left(1\right)$ corecursively as
$\sigma\left(\chi\right)=\left(\sum_{s\in S}\sum_{a\in A}\mathbb{P}_{\chi}\left(s,\top,a\right),\right.\\ \left.\left(\sum_{s\in S}\sum_{a\in A}\mathbb{P}_{\chi}\left(s,\bot,a\right)\right)\sigma\left(\sum_{s\in S}\sum_{a\in A}\frac{\mathbb{P}_{\chi}\left(s,\bot,a\right)}{\sum_{s^{\prime}\in S}\sum_{a^{\prime}\in A}\mathbb{P}_{\chi}\left(s^{\prime},\bot,a^{\prime}\right)}\mathbb{T}{\chi}\left(\star,\left(s,\bot,a\right)\right)\right)\right).$ (53)
One should think of this as a sequence of probabilities of the form
$p(\text{success occurs for the first time on time step $n$}).$ (54)
Definition 4.2.2 lets us extend the notion of a mixture to bounded sum sequences, suggesting the following lemma.
Lemma 4.2.4 (Success sequences respect mixtures).
Let $\chi_{k}\colon\{\star\}\mathrel{\triangleright}S\times G\times A$ be coupled transducers and $\alpha_{k}\geq 0$, such that $\sum_{k=0}^{\infty}\alpha_{k}=1$. Success sequences respect mixtures, that is
$\sigma\left(\sum_{k=1}^{\infty}\alpha_{k}\chi_{k}\right)=\sum_{k=1}^{\infty}\alpha_{k}\sigma\left(\chi_{k}\right).$ (55)
Proof.
The first projection of the left hand side (i.e. the first element of the pair returned by $\sigma$) is
$\sum_{s\in S}\sum_{a\in A}\mathbb{P}_{\sum_{k=1}^{\infty}\alpha_{k}\chi_{k}}\left(s,\top,a\right)=\sum_{s\in S}\sum_{a\in A}\sum_{k=1}^{\infty}\alpha_{k}\mathbb{P}_{\chi_{k}}\left(s,\top,a\right),$ (56)
which after simple operations on sums becomes
$\sum_{k=1}^{\infty}\alpha_{k}\sum_{s\in S}\sum_{a\in A}\mathbb{P}_{\chi_{k}}\left(s,\top,a\right),$ (57)
which is exactly the first projection of the right hand side.
The second projection is the slightly more complex
$\left(\sum_{s\in S}\sum_{a\in A}\sum_{k=0}^{\infty}\alpha_{k}\mathbb{P}_{\chi_{k}}\left(s,\bot,a\right)\right)\\ \sigma\left(\sum_{s\in S}\sum_{a\in A}\frac{\sum_{k=0}^{\infty}\alpha_{k}\mathbb{P}_{\chi_{k}}\left(s,\bot,a\right)}{\sum_{s^{\prime}\in S}\sum_{a^{\prime}\in A}\sum_{k=0}^{\infty}\alpha_{k}\mathbb{P}_{\chi_{k}}\left(s^{\prime},\bot,a^{\prime}\right)}\right.\\ \left.\sum_{k=0}^{\infty}\frac{\alpha_{k}\mathbb{P}_{\chi_{k}}\left(s,\bot,a\right)}{\sum_{k^{\prime}=0}^{\infty}\alpha_{k}^{\prime}\mathbb{P}_{\chi_{k^{\prime}}}\left(s,\bot,a\right)}\mathbb{T}{\chi_{k}}\left(\star,\left(s,\bot,a\right)\right)\right).$ (58)
Canceling out the numerator of the first fraction with the denominator of the second we get
$\left(\sum_{s\in S}\sum_{a\in A}\sum_{k=0}^{\infty}\alpha_{k}\mathbb{P}_{\chi_{k}}\left(s,\bot,a\right)\right)\\ \sigma\left(\sum_{s\in S}\sum_{a\in A}\sum_{k=0}^{\infty}\frac{\alpha_{k}\mathbb{P}_{\chi_{k}}\left(s,\bot,a\right)}{\sum_{s^{\prime}\in S}\sum_{a^{\prime}\in A}\sum_{k^{\prime}=0}^{\infty}\alpha_{k^{\prime}}\mathbb{P}_{\chi_{k^{\prime}}}\left(s^{\prime},\bot,a^{\prime}\right)}\mathbb{T}{\chi_{k}}\left(\star\left(s,\bot,a\right)\right)\right).$ (59)
Expanding $1$ into a fraction with ${\sum_{s^{\prime}\in S}\sum_{a^{\prime}\in A}\mathbb{P}_{\chi_{k}}\left(s^{\prime},\bot,a^{\prime}\right)}$ for every $k$ inside of the corecursive function and shuffling around some terms yields
$\left(\sum_{s\in S}\sum_{a\in A}\sum_{k=0}^{\infty}\alpha_{k}\mathbb{P}_{\chi_{k}}\left(s,\bot,a\right)\right)\\ \sigma\left(\sum_{k=0}^{\infty}\frac{\sum_{s^{\prime}\in S}\sum_{a^{\prime}\in A}\alpha_{k}\mathbb{P}_{\chi_{k}}\left(s^{\prime},\bot,a^{\prime}\right)}{\sum_{s^{\prime}\in S}\sum_{a^{\prime}\in A}\sum_{k^{\prime}=0}^{\infty}\alpha_{k^{\prime}}\mathbb{P}_{\chi_{k^{\prime}}}\left(s^{\prime},\bot,a^{\prime}\right)}\right.\\ \left.\sum_{s\in S}\sum_{a\in A}\frac{\mathbb{P}_{\chi_{k}}\left(s,\bot,a\right)}{\sum_{s^{\prime}\in S}\sum_{a^{\prime}\in A}\mathbb{P}_{\chi_{k}}\left(s^{\prime},\bot,a^{\prime}\right)}\mathbb{T}{\chi_{k}}\left(\star,\left(s,\bot,a\right)\right)\right),$ (60)
and finally using the coinductive hypothesis, cancelling the first term with the first denominator, and reorganizing some sums we get
$\sum_{k=0}^{\infty}\alpha_{k}\left(\sum_{s\in S}\sum_{a\in A}\mathbb{P}_{\chi_{k}}\left(s,\bot,a\right)\right)\\ \sigma\left(\sum_{s\in S}\sum_{a\in A}\frac{\mathbb{P}_{\chi_{k}}\left(s,\bot,a\right)}{\sum_{s^{\prime}\in S}\sum_{a^{\prime}\in A}\mathbb{P}_{\chi_{k}}\left(s^{\prime},\bot,a^{\prime}\right)}\mathbb{T}{\chi_{k}}\left(\star,\left(s,\bot,a\right)\right)\right),$ (61)
which is exactly the second term of the right hand side, finishing the proof. ∎
This lets us define the probability of success. We start by defining the sum of a BSS:
Definition 4.2.5 (Sum of a BSS).
Given a BSS, we define its $n$-step sum
$\Sigma_{n}\colon\mathrm{BSS}\left(r_{0}\right)\to\left[0,r_{0}\right]$ (62)
inductively with the base
$\Sigma_{0}\left(\left(r,t\right)\right)=r$ (63)
and the inductive step
$\Sigma_{n+1}\left(\left(r,t\right)\right)=r+\Sigma_{n}\left(t\right).$ (64)
A simple inductive argument shows that this is well defined.
The sequence $\left(\Sigma_{n}\left(\left(r,t\right)\right)\right)_{n\in\mathbb{N}}$ is obviously bounded and another simple inductive argument shows that it is non-decreasing, letting us define the sum
$\Sigma\left(\left(r,t\right)\right)=\lim_{n\to\infty}\Sigma_{n}\left(\left(r,t\right)\right).$ (65)
This, again, respects mixtures.
Lemma 4.2.6 (Sum of BSSes respect mixtures).
Let $t_{k}\colon\mathrm{BSS}\left(r_{0,k}\right)$ be BSSes and $\alpha_{k}\geq 0$, such that $\sum_{k=0}^{\infty}\alpha_{k}=1$. The sum of BSSes respects mixtures, that is
$\Sigma\left(\sum_{k=1}^{\infty}\alpha_{k}t_{k}\right)=\sum_{k=1}^{\infty}\alpha_{k}\Sigma\left(t_{k}\right)$ (66)
Proof.
It suffices to note that converging limits respect mixtures and perform a simple inductive proof
$\Sigma_{0}\left(\sum_{k=1}^{\infty}\alpha_{k}\left(r_{k},t_{k}\right)\right)=\sum_{k=1}^{\infty}\alpha_{k}r_{k}=\sum_{k=1}^{\infty}\alpha_{k}\Sigma_{0}\left(\left(r_{k},t_{k}\right)\right)$ (67)
with the step
$\Sigma_{n+1}\left(\sum_{k=1}^{\infty}\alpha_{k}\left(r_{k},t_{k}\right)\right)=\sum_{k=1}^{\infty}\alpha_{k}r_{k}+\Sigma_{n}\left(\sum_{k=1}^{\infty}\alpha_{k}t_{k}\right)=\\ \sum_{k=1}^{\infty}\alpha_{k}r_{k}+\sum_{k=1}^{\infty}\alpha_{k}\Sigma_{n}\left(t_{k}\right)=\sum_{k=1}^{\infty}\alpha_{k}\Sigma_{n+1}\left(\left(r_{k},t_{k}\right)\right).$ (68)
With this, we can define the success probability:
Definition 4.2.7.
We define the success probability of a policy-environment pair as the composition
$S\left(\pi,\varepsilon\right)=\Sigma\left(\sigma\left(c\left(\pi,\varepsilon\right)\right)\right)$ (69)
and analogously for $n$-step success probabilities.
Because of the form of the inductive definition success probability satisfies the following equation:
$S\left(\pi,\varepsilon\right)=\sum_{s\in S}\sum_{a\in A}\mathbb{P}_{c\left(\pi,\varepsilon\right)}\left(s,\top,a\right)+\mathbb{P}_{c\left(\pi,\varepsilon\right)}\left(s,\bot,a\right)S\left(\mathbb{T}_{c\left(\pi,\varepsilon\right)}\left(\star,\left(s,\bot,a\right)\right)\right).$ (70)
Intuitively the first term in the definition of probability of success represents the probability of succeeding in a single step, while the second term corresponds to succeeding at any point in the future, assuming no success has been achieved in this step.
One could be tempted to use equation (70) as a ‘corecursive’ definition of success probability, but that would not work, because the codomain $\left[0,1\right]$ is not coinductively defined by $S$. Indeed, this equation is also satisfied by the constant function $1$.
To make the meaning of equation (70) clearer consider the following equations for the $n$-step probability of success of a specific policy-environment pair.
$S_{0}\left(\pi,\varepsilon\right)$ $=\sum_{s\in S}\sum_{a\in A}\mathbb{P}_{c\left(\pi,\varepsilon\right)}\left(s,\top,a\right),$
$S_{n+1}\left(\pi,\varepsilon\right)$ $=\sum_{s\in S}\sum_{a\in A}\mathbb{P}_{c\left(\pi,\varepsilon\right)}\left(s,\top,a\right)+\mathbb{P}_{c\left(\pi,\varepsilon\right)}\left(s,\bot,a\right)S_{n}\left(\mathbb{T}_{c\left(\pi,\varepsilon\right)}\left(\star,\left(s,\bot,a\right)\right)\right).$
For a given evolution $\left(\mathbf{s}:\mathbf{a}\right)\in\left(S\times A\right)^{n}$ we can define the failure probability along this evolution
$F_{c\left(\pi,\varepsilon\right)}\left(\mathbf{s},\mathbf{a}\right)=\begin{cases}\prod_{i=0}^{n}\mathbb{P}_{as(i)}\left(\mathbf{s}_{i+1},\bot\right)&\textrm{ if }\pi\bullet\left(\mathbf{s}:\mathbf{a}\right)\textrm{ is valid,}\\ 0&\textrm{ otherwise.}\end{cases}$ (71)
The subscript $as(i)$ is defined as $\varepsilon\bullet\left(\mathbf{a}_{\leq i}:\left(\mathbf{s}_{\leq i}:\bot^{i}\right)\right)$ and represents the evolution of the environment following $\left(\mathbf{s}:\mathbf{a}\right)$ for $i$ steps and assuming no success at every step.
Note that this function is zero if the environment evolution is not valid, because one of the elements of the product will be zero.
Now, by expanding equation (70) $n$ times, we can write down an alternative formulation of the success probability for each $n$
$S\left(\pi,\varepsilon\right)=S_{n}\left(\pi,\varepsilon\right)+\sum_{\mathbf{s}\in S^{n}}\sum_{\mathbf{a}\in A^{n}}F_{c\left(\pi,\varepsilon\right)}\left(\mathbf{s},\mathbf{a}\right)S\left(\pi\bullet\left(\mathbf{s}:\mathbf{a}\right),\varepsilon\bullet\left(\mathbf{a}:\left(\mathbf{s}:\bot^{n}\right)\right)\right),$ (72)
which is a sum of the probability of success within the first $n$ steps, plus the expected value of the probability of success after a valid $n$-step evolution which did not achieve success.
Probability of success also has the very useful property of respecting mixtures both in policies as well as environments.
Lemma 4.2.8 (Success probabilities respect mixtures).
Let $\pi_{k}\colon S\mathrel{\triangleright}A$ be policies, $\varepsilon_{k}\colon A\mathrel{\triangleright}\left(S\times G\right)$ be environments, and $\alpha_{k},\beta_{k}\geq 0$ such that $\sum_{k=0}^{\infty}\alpha_{k}=1$ and $\sum_{k=0}^{\infty}\beta_{k}=1$. The success probability function respects mixtures, that is
$S\left(\sum_{k=1}^{\infty}\alpha_{k}\pi_{k},\sum_{k=1}^{\infty}\beta_{k}\varepsilon_{k}\right)=\sum_{k=1}^{\infty}\sum_{l=1}^{\infty}\alpha_{k}\beta_{l}S\left(\pi_{k},\varepsilon_{l}\right)$ (73)
Proof.
By Lemma 4.1.2 coupling respects mixtures, by Lemma 4.2.4 success sequences respect mixtures, and by Lemma 4.2.6 sums of BSSes respect mixtures, thus so does their composition. ∎
With success defined and explored we can now define what it means for a policy to be optimal for an environment.
Definition 4.2.9 (Optimality).
Let $T\subseteq S\mathrel{\triangleright}A$ be a constrained set of policies. We will say that a policy $\pi\colon T$ is $T$-optimal for a teleo-environment $\varepsilon\colon A\mathrel{\triangleright}\left(S\times G\right)$ iff
$\forall_{\pi^{\prime}\colon T}S\left(\pi^{\prime},\varepsilon\right)\leq S\left(\pi,\varepsilon\right).$ (74)
In other words $\pi$ maximizes the success probability in $\varepsilon$ among the policies in $T$. We will usually omit the ‘$T$-’, since it will be obvious from context.
5 Filtering
We remarked in Section 4.1.1 that the evolution operator can be seen in terms of Bayesian filtering. We will now explain this intuition in more detail, and derive some formal results.
An optimal policy $\pi$ for $\varepsilon$ represents the behavioural propensities of an ideally-performing agent with beliefs and values encoded by $\varepsilon$. In this sense, it is consistent with $\pi$ to attribute a ‘mental state’ represented by $\varepsilon$ (under the assumption that $\pi$ behaves ‘rationally’). We can say, as it were, that $\varepsilon$ is a ‘permissible’ mental state attribution for $\pi$, at least for practical purposes.
Iterated application of the evolution operator for a policy $\pi$ produces a sequence of transducers $\pi_{1},\pi_{2},\cdots$, and so on, where $\pi_{n+1}=\pi_{n}\bullet(s_{n},a_{n})$ for some $(s_{n},a_{n})$. This section will establish that, if it is permissible to attribute a ‘mental state’ $\varepsilon_{1}$ to $\pi_{1}$, then it is permissible to attribute some $\varepsilon_{n}$ to each $\pi_{n}$ so that the sequence $\varepsilon_{1},\varepsilon_{2},\cdots$ is what would be obtained by a process resembling Bayesian filtering, beginning with a prior corresponding to $\varepsilon_{1}$, and proceeding at each step by updating on an observation corresponding to $(a_{n},(s_{n},\bot))$.
It is worth pointing out that this differs from standard Bayesian filtering in that it is ‘value-laden’: according to the sequence $\varepsilon_{1},\varepsilon_{2},\cdots$, the transition $\pi_{n}\to\pi_{n+1}$ ‘behaves as though’ an agent were updating on a notional observation $\bot$ over the environment’s ‘telos’ channel, in addition to observations of $\pi_{n}$’s actual input $s_{n}$ and output $a_{n}$. Recall that $\pi_{n}$ does not take a telos signal as an input, although $\varepsilon_{n}$ produces it as an output. We will see in Section 5.2 that if there exists any $\varepsilon$ for which $\pi_{1}$ is optimal, it is also admissible to attribute a sequence $\varepsilon_{1},\varepsilon_{2},\cdots$ according to which $\pi_{n}\to\pi_{n+1}$ ‘behaves as though’ the agent were updating only on the actual values $s_{n}$ and $a_{n}$ (and not an imaginary observation $\bot$).
The reason for the value-laden nature of this filtering can be explained as follows. The way to achieve the best chances of future success in a new environment may depend on whether or not the current environment emitted the success signal. But the agent’s goal is to achieve success at least once, with success signals after that counting for nothing. If success has been attained in the current time step, there is consequently no advantage to pursuing an optimal policy in the next step – if the agent doesn’t know whether its goal was attained, it shouldn’t be distracted by the possibility that it was. Hence, the agent should always behave as though it has not already achieved success, in order to maximise its overall chances of eventual success. We will discuss this more in Section 6.6.
Our formal treatment considers (value-laden) filtering for the more general case of constrained policies (i.e. ones that are $T$-optimal for some set $T$) rather than only policies that are globally optimal. We will begin by defining something we call the ‘value-laden Bellman property’.
Definition 5.0.1 (Value-laden Bellman property for transducers).
Let $T\subseteq S\mathrel{\triangleright}A$ be a set of constrained policies. We say that $T$ has the Bellman property for transducers if for any policy $\pi\colon T$ that is $T$-optimal for an environment $\varepsilon\colon A\mathrel{\triangleright}\left(S\times G\right)$ and any trajectory $\left(\mathbf{a},\mathbf{s}\right)\in A^{n}\times S^{n}$ of length $n$, if $\varepsilon^{\prime}=\varepsilon\bullet\left(\mathbf{a}:\left(\mathbf{s}:\bot^{n}\right)\right)$ and $\pi^{\prime}=\pi\bullet\left(\mathbf{s}:\mathbf{a}\right)$ are valid, then $\pi^{\prime}$ is $T$-optimal for $\varepsilon^{\prime}$.
This property says that if an agent is optimal within its class for one problem (in the sense of maximising success probability), then after it has interacted with the environment for some time it should still be optimal in its new, resulting environment. This is essentially Bellman’s principle of optimality [4, Chapter III, §3]. However, a difference in our case is that to obtain the environment for the next time step we condition not only on the agent’s action and received sensor value, but also on success not occurring in the current time step. As noted above, this is because the agent’s goal is to achieve success at least once. The agent does not know whether success has been achieved or not, but its future actions will only matter in the case where success has not already been achieved. We will say more on this point below.
We want to know which sets of constrained policies have the Bellman property. We can prove the following theorem.
Theorem 5.0.2 (Value-laden Bellman theorem for transducers).
Let $T\subseteq S\mathrel{\triangleright}A$ be a set of constrained policies. If $T$ has the exchange property from Definition 3.3.1, it has the Bellman property.
Corollary 5.0.3.
The set of all transducers $S\mathrel{\triangleright}A$ has the Bellman property.
Proof of Theorem 5.0.2.
We will prove this by induction on the length of the trajectory. For $n=0$ the theorem is trivially true, because $\varepsilon^{\prime}=\varepsilon$ and $\pi^{\prime}=\pi$. Let us then assume that the theorem is true for $n-1$.
Let us write $\pi^{\prime\prime}=\pi\bullet\left(\mathbf{s}_{<n}:\mathbf{a}_{<n}\right)$, and $\varepsilon^{\prime\prime}=\varepsilon\bullet\left(\mathbf{a}_{<n}:\left(\mathbf{s}_{<n}:\bot^{n-1}\right)\right)$. Then, by induction, $\pi^{\prime\prime}$ is optimal for $\varepsilon^{\prime\prime}$. Since $\pi^{\prime}=\pi^{\prime\prime}\bullet\left(\mathbf{s}_{n}:\mathbf{a}_{n}\right)$ and $\varepsilon^{\prime}=\varepsilon^{\prime\prime}\bullet\left(\mathbf{a}_{n}:\left(\mathbf{s}_{n}:\bot\right)\right)$ it is then sufficient to prove that the theorem holds for evolution by a single action-state pair $\left(a,s\right)$.
Assume then that $\pi^{\prime}=\pi\bullet\left(s,a\right)=\mathbb{T}_{\pi}\left(s,a\right)$ is not optimal for $\varepsilon^{\prime}=\varepsilon\bullet\left(a,\left(s,\bot\right)\right)$, that is there exists a policy $\pi^{*^{\prime}}$, such that $S\left(\pi^{*^{\prime}},\varepsilon^{\prime}\right)>S\left(\mathbb{T}_{\pi}\left(s,a\right),\varepsilon^{\prime}\right)$. We will construct a policy $\pi^{*}\colon T$ such that $S\left(\pi^{*},\varepsilon\right)>S\left(\pi,\varepsilon\right)$, thus contradicting the fact that $\pi$ is optimal for $\varepsilon$.
Define this policy as $\pi^{*}=\left(\mathbb{P}_{\pi},t\right)$, where
$t\left(i,o\right)=\begin{cases}\pi^{*^{\prime}}&\textrm{ if }\left(i,o\right)=\left(s,a\right),\\ \mathbb{T}_{\pi}\left(i,o\right)&\textrm{ otherwise.}\end{cases}$ (75)
This belongs to $T$ since $T$ has the exchange property.
Now it remains to show that $S\left(\pi^{*},\varepsilon\right)<S\left(\pi,\varepsilon\right).$ Using equation (70) considering the form of $\pi^{*}$ we get
$S\left(\pi^{*},\varepsilon\right)=\sum_{u\in S}\sum_{b\in A}\mathbb{P}_{c\left(\pi,\varepsilon\right)}\left(u,\top,b\right)+\mathbb{P}_{c\left(\pi,\varepsilon\right)}\left(u,\bot,b\right)S\left(t\left(u,b\right),\mathbb{T}_{\varepsilon}\left(b,\left(u,\bot\right)\right)\right).$ (76)
Expanding the definition of $t$ and separating the term for $\left(a,s\right)$ we get
$S\left(\pi^{*},\varepsilon\right)=\sum_{u\in S}\sum_{\begin{subarray}{c}b\in A\\ \left(b,u\right)\neq\left(a,s\right)\end{subarray}}\left(\mathbb{P}_{c\left(\pi,\varepsilon\right)}\left(u,\top,b\right)+\mathbb{P}_{c\left(\pi,\varepsilon\right)}\left(u,\bot,b\right)S\left(\mathbb{T}_{\pi}\left(u,b\right),\mathbb{T}_{\varepsilon}\left(b,\left(u,\bot\right)\right)\right)\right)\\ +\mathbb{P}_{c\left(\pi,\varepsilon\right)}\left(s,\top,a\right)+\mathbb{P}_{c\left(\pi,\varepsilon\right)}\left(s,\bot,a\right)S\left(\pi^{*^{\prime}},\mathbb{T}_{\varepsilon}\left(a,\left(s,\bot\right)\right)\right).$ (77)
Noting that $\mathbb{T}_{\varepsilon}\left(a,\left(s,\bot\right)\right)=\varepsilon^{\prime}$ and using the inequality on probabilities of success considering that $\mathbb{P}_{c\left(\pi,\varepsilon\right)}\left(s,\bot,a\right)\neq 0$ because the evolution of the environment was valid, we can now write
$S\left(\pi^{*},\varepsilon\right)>\sum_{u\in S}\sum_{\begin{subarray}{c}b\in A\\ \left(b,u\right)\neq\left(a,s\right)\end{subarray}}\left(\mathbb{P}_{c\left(\pi,\varepsilon\right)}\left(u,\top,b\right)+\mathbb{P}_{c\left(\pi,\varepsilon\right)}\left(u,\bot,b\right)S\left(\mathbb{T}_{\pi}\left(u,b\right),\mathbb{T}_{\varepsilon}\left(b,\left(u,\bot\right)\right)\right)\right)\\ +\mathbb{P}_{c\left(\pi,\varepsilon\right)}\left(s,\top,a\right)+\mathbb{P}_{c\left(\pi,\varepsilon\right)}\left(s,\bot,a\right)S\left(\mathbb{T}_{\pi}\left(s,a\right),\mathbb{T}_{\varepsilon}\left(a,\left(s,\bot\right)\right)\right),$ (78)
and after reintegrating the last term we can use equation (70) to get $S\left(\pi^{*},\varepsilon\right)>S\left(\pi,\varepsilon\right),$ finishing the proof. ∎
While satisfying the exchange property is a sufficient condition for the Bellman property to hold it is not necessary. Consider the one-flip transducers defined in Example 3.3.5. They do not have the exchange property, but the Bellman theorem still holds for them. This is due to the fact that, as we will see later, for any environment there is always an optimal deterministic policy and a ‘limited exchange property’, where you only replace future policies with deterministic ones, holds for one-flip transducers.
The fact that transducers obey the (value-laden) Bellman theorem has an important consequence. Let us take stock of what it entails.
Our thesis is that if a physical system is optimal for some problem (specified, in our case, as a teleo-environment), then we can attribute beliefs and goals to the system that correspond to the given problem. In our case, this means that if a policy $\pi$ is optimal for a given teleo-environment $\varepsilon$, then it is permissible to attribute to $\pi$ the goal of achieving the telos signal at least once, and the belief that the dynamics of the environment (and the telos signal) are given by $\varepsilon$. Corollary 5.0.3 says that attributing beliefs in this way is consistent with Bayesian updating in the following sense:
Suppose we can attribute beliefs $\varepsilon$ to $\pi$, and that $\pi$ then emits action $a:A$ and receives sensor value $s:S$ from the environment, evolving into $\pi\bullet(s,a)$. It is then permissible to attribute beliefs $\varepsilon\bullet(a,s,\bot)$ to $\pi\bullet(s,a)$. We have established in Section 3 that evolving a transducer is closely related to Bayesian filtering. Thus, after the update, $\pi\bullet(s,a)$ can be attributed beliefs that corresponding to performing a Bayesian filtering update on the environment.
In this story the filtering step conditions not only on the agent’s received sensor data $s$ but also on success not being achieved on the current time step (i.e. the $\bot$ symbol), even though the agent doesn’t have access to the telos channel, and indeed success might actually have occurred. We argued that we can understand this ‘from the agent’s point of view’ by noting that since the agent’s (attributed) goal is to achieve success at least once, if success has already occurred, the agent’s actions no longer matter. As we will eventually see, in Example 6.6.2, if an agent would condition only on $s$ and $a$ but not $\bot$, it would not necessarily be optimal for the resulting conditioned environment.
5.1 UFS transducers do not filter
There are classes of transducers for which the Bellman property does not hold. For a fairly general and nontrivial example we will look at unifilar finite state transducers. First we will require some definitions, which will also be used for future examples and proofs.
Definition 5.1.1 (Success and nothing distributions).
Let $S$ be an $n$-element state space. We define the (uniform) success and nothing distributions on $S\times G$ as
$U_{\top}\left(s,t\right)=\begin{cases}\frac{1}{n}&\textrm{ if }t=\top,\\ 0&\textrm{ otherwise.}\end{cases}$ (79)
and
$U_{\bot}\left(s,t\right)=\begin{cases}\frac{1}{n}&\textrm{ if }t=\bot,\\ 0&\textrm{ otherwise.}\end{cases}$ (80)
respectively.
Definition 5.1.2 (Doom and despair environments).
The doom environment is one that never returns success, formally defined as
$\varepsilon_{0}=\left(U_{\bot},\left(a,\left(s,t\right)\right)\mapsto\varepsilon_{0}\right),$ (81)
while the despair environment has $\frac{1}{2}$ probability of success for one step and then changes into the doom environment, regardless of which action is taken
$\varepsilon_{\frac{1}{2}}=\left(\frac{1}{2}U_{\bot}+\frac{1}{2}U_{\top},\left(a,\left(s,t\right)\right)\mapsto\varepsilon_{0}\right).$ (82)
Example 5.1.3 (No UFS filtering).
Let $S=\left\{1,\dots,n\right\}$, $A=\left\{1,\dots,n,n+1\right\}$ and consider only policies constrained to $\mathrm{UFS}_{n}\subseteq S\mathrel{\triangleright}A$. We will build an explicit counterexample to filtering in this setting, defining the necessary environment step by step.
First we define a family of (weighted) mimic environments indexed by states
$\varepsilon_{m}\left(s^{\prime}\right)=\left(\frac{2^{s^{\prime}}-1}{2^{s^{\prime}}}U_{\bot}+\frac{1}{2^{s^{\prime}}}U_{\top},\left(a,\left(s,t\right)\right)\mapsto\begin{cases}\varepsilon_{m}\left(s\right)&\textrm{ if }a=s^{\prime},\\ \varepsilon_{\frac{1}{2}}&\textrm{ if }a=n+1,\\ \varepsilon_{0}&\textrm{ otherwise}\end{cases}\right).$ (83)
The name refers to the fact that an optimal policy will have to mimic what the environment is doing: if the previous sensor input was $k$ then the next action should also be $k$.
It now remains to define the environment itself
$\varepsilon=\left(U_{\bot},\left(a,\left(s,t\right)\right)\mapsto\begin{cases}\varepsilon_{m}\left(s\right)&\textrm{ if }a=n+1,\\ \varepsilon_{0}&\textrm{ otherwise}\end{cases}\right).$ (84)
In other words, the environment never returns success on the first time step, and the action must be $n+1$.
The policy induced by the unifilar $n$ state machine (with $X=A$)
$\pi^{*^{\prime}}=\left(s^{\prime},x\mapsto x,\left(x,s,a\right)\mapsto s\right)$ (85)
is optimal for every mimic environment $\varepsilon_{m}\left(s^{\prime}\right)$. (We abuse notation and write $s^{\prime}$ for the deterministic distribution always returning $s^{\prime}$.) This policy has success probability $1$, since it can succeed with constant expected probability in every step, so it will eventually succeed. Any other policy will have a nonzero probability of not repeating the previous state as its action, which would make it fail forever (since then the environment gives at most one constant probability of success and afterwards behaves like the doom environment), giving a probability of success below $1$.
The policy induced by the unifilar $n$ state machine
$\pi=\left(n,x\mapsto\begin{cases}x&\textrm{ if }x\neq n,\\ n+1&\textrm{ if }x=n\end{cases},\left(x,s,a\right)\mapsto s\right).$ (86)
is optimal for the environment itself.
This transducer is (uniquely) optimal for $\varepsilon$ within the class of $n$-state UFS transducers but it is not optimal among all transducers. (In fact there is an $(n+1)$-state UFS transducer that outperforms it.)
We argue that $\pi$ is optimal informally rather than giving a proof. The policy has to return $n+1$ as its first action, otherwise its probability of success is 00. This constrains the output function to map some memory state to $n+1$ making it impossible for the future function to be a perfect mimic. Mimicking the first $n-1$ states/actions is the uniquely optimal solution, because they have the best chance of providing success when they occur, while all states have the same chance of occurring. At the same time when we no longer can mimic the environment, the policy is optimal when it returns $n+1$ for another shot at a success. In the end the success probability of such an imperfect mimic is $\frac{2^{n-1}+2^{n-2}-1}{2^{n}-1}$ and any other unifilar $n$ state will have a lower result, although it is tedious to check.
Any evolution of $\pi$ has a success probability lower than evolutions of $\pi^{*^{\prime}}$ (as it is lower than $1$), contradicting filtering.
Interestingly the environment in the above example can itself be defined as a UFS machine with $n+3$ states – one for the initial state, one for despair, one for doom, and $n$ for the usual states remembering history. We do not know whether the constant $3$ can be lowered.
5.2 Sensorimotor-only Filtering
Theorem 5.0.2 requires that the environment is evolved under the assumption that no success is returned at every step. We will explore how important this assumption really is, and in which situations it can be dropped. First let us define what dropping this assumption would even look like.
Definition 5.2.1 (Success-ambivalent evolution).
Let $\varepsilon\colon A\mathrel{\triangleright}\left(S\times G\right)$ be an environment. We define its success-ambivalent evolution to be
$\varepsilon\bullet\left(a,s\right)=\frac{\mathbb{P}_{\varepsilon}\left(\left(s,\bot\right)\right)}{\mathbb{P}_{\varepsilon}\left(\left(s,\bot\right)\right)+\mathbb{P}_{\varepsilon}\left(\left(s,\top\right)\right)}\mathbb{T}_{\varepsilon}\left(a,\left(s,\bot\right)\right)+\\ \frac{\mathbb{P}_{\varepsilon}\left(\left(s,\top\right)\right)}{\mathbb{P}_{\varepsilon}\left(\left(s,\bot\right)\right)+\mathbb{P}_{\varepsilon}\left(\left(s,\top\right)\right)}\mathbb{T}_{\varepsilon}\left(a,\left(s,\top\right)\right),$ (87)
that is the mixture of transducers resulting from either a success or no success transition proportional to the relative probabilities of such transitions.
We use the same notation as for normal evolutions, but context should disambiguate sufficiently well to avoid confusing the reader.
Now consider a special class of environments that achieve success at most once at a given trajectory.
Definition 5.2.2 (Single success environments).
Let $\varepsilon\colon A\mathrel{\triangleright}\left(S\times G\right)$ be an environment. We call it a single success environment if no trajectory $\left(\mathbf{a}:\left(\mathbf{s}:\mathbf{t}\right)\right)$ such that there exist two indices $i,j\in\mathbb{N}$ for which $\mathbf{t}_{i}=\mathbf{t}_{j}=\top$ is a valid evolution for $\varepsilon$. We denote the set of all such environments as $W$.
We can inject any environment into $W$.
Definition 5.2.3 (Single success truncation).
We define the single success truncation to be a function $Z\colon A\mathrel{\triangleright}\left(S\times G\right)\to W$ defined as
$Z\left(\varepsilon\right)=\left(\mathbb{P}_{\varepsilon},\left(a,\left(s,t\right)\right)\mapsto\begin{cases}Z\left(\mathbb{T}_{\varepsilon}\left(a,\left(s,t\right)\right)\right)&\textrm{ if }t=\bot,\\ D\left(\mathbb{T}_{\varepsilon}\left(a,\left(s,t\right)\right)\right)&\textrm{if }t=\top\end{cases}\right),$ (88)
where $D\colon A\mathrel{\triangleright}\left(S\times G\right)\to A\mathrel{\triangleright}\left(S\times G\right)$ is the dooming function that removes all successes, defined as
$D\left(\varepsilon\right)=\left(d\left(\mathbb{P}_{\varepsilon}\right),\left(a,\left(s,t\right)\right)\mapsto D\left(\varepsilon\bullet\left(a,s\right)\right)\right),$ (89)
with $d$ defined simply as
$d\left(p\right)\left(\left(s,t\right)\right)=\begin{cases}p\left(\left(s,\top\right)\right)+p\left(\left(s,\bot\right)\right)&\textrm{ if }t=\bot,\\ 0&\textrm{ otherwise.}\end{cases}$ (90)
It’s easy to see that $Z$ is an identity on $W$, justifying the name.
It turns out that, because we only care about success ever occurring, single success truncation preserves success probability.
Lemma 5.2.4 (Truncation preserves success probability).
Let $\pi\colon S\mathrel{\triangleright}A$ and $\varepsilon\colon A\mathrel{\triangleright}\left(S\times G\right)$ be an arbitrary policy-environment pair. Then
$S\left(\pi,\varepsilon\right)=S\left(\pi,Z\left(\varepsilon\right)\right).$ (91)
Proof.
We will prove this for success sequences and apply Lemma 6.4.3.
Since truncation preserves associated probabilities, it sufficess to inspect the transition function
$\sum_{s\in S}\sum_{a\in A}\mathbb{P}_{c\left(\pi,Z\left(\varepsilon\right)\right)}\left(s,\bot,a\right)\sigma\left(c\left(\mathbb{T}_{\pi}\left(s,a\right),\mathbb{T}_{Z\left(\varepsilon\right)}\left(a,\left(s,\bot\right)\right)\right)\right)=$ (92)
applying the definition of single success truncation
$\sum_{s\in S}\sum_{a\in A}\mathbb{P}_{c\left(\pi,\varepsilon\right)}\left(s,\bot,a\right)\sigma\left(c\left(\mathbb{T}_{\pi}\left(s,a\right),Z\left(\mathbb{T}_{\varepsilon}\left(a,\left(s,\bot\right)\right)\right)\right)\right).$ (93)
It remains to use the coinductive hypothesis to finish the proof. ∎
In particular this means that if a policy is optimal for an environment $\varepsilon$, then it is also optimal for its truncated version $Z\left(\varepsilon\right)$.
There is one more lemma that will prove useful in the upcoming theorem proof.
Lemma 5.2.5.
Let $\pi\colon S\mathrel{\triangleright}A$ be a policy, $\varepsilon\colon W$ be a single success environment, and $\left(a,s\right)\in A\times S$ be an action-state pair. Then
$S\left(\pi,\varepsilon\bullet\left(a,s\right)\right)=\frac{\mathbb{P}_{\varepsilon}\left(\left(s,\bot\right)\right)}{\mathbb{P}_{\varepsilon}\left(\left(s,\bot\right)\right)+\mathbb{P}_{\varepsilon}\left(\left(s,\top\right)\right)}S\left(\pi,\varepsilon\bullet\left(a,\left(s,\bot\right)\right)\right).$ (94)
Proof.
By the definition of success-ambivalent evolution and Lemma 4.2.8 we can write the left hand side as
$\frac{\mathbb{P}_{\varepsilon}\left(\left(s,\bot\right)\right)}{\mathbb{P}_{\varepsilon}\left(\left(s,\bot\right)\right)+\mathbb{P}_{\varepsilon}\left(\left(s,\top\right)\right)}S\left(\pi,\varepsilon\bullet\left(a,\left(s,\bot\right)\right)\right)+\\ \frac{\mathbb{P}_{\varepsilon}\left(\left(s,\top\right)\right)}{\mathbb{P}_{\varepsilon}\left(\left(s,\bot\right)\right)+\mathbb{P}_{\varepsilon}\left(\left(s,\top\right)\right)}S\left(\varepsilon\bullet\left(a,\left(s,\top\right)\right)\right),$ (95)
but since $\varepsilon$ is single success, $S\left(\varepsilon\bullet\left(a,\left(s,\top\right)\right)\right)=0$, leaving
$\frac{\mathbb{P}_{\varepsilon}\left(\left(s,\bot\right)\right)}{\mathbb{P}_{\varepsilon}\left(\left(s,\bot\right)\right)+\mathbb{P}_{\varepsilon}\left(\left(s,\top\right)\right)}S\left(\pi,\varepsilon\bullet\left(a,\left(s,\bot\right)\right)\right),$ (96)
as required. ∎
This allows us to prove the sensorimotor-only filtering result.
Theorem 5.2.6 (Sensorimotor-only Bellman theorem for transducers).
Let $T\subseteq S\mathrel{\triangleright}A$ be a set of constrained policies, which satisfies the exchange property. Let $\pi\colon T$ be a policy optimal for the single success environment $\varepsilon\colon W$ and $\left(\mathbf{a},\mathbf{s}\right)\in A^{n}\times S^{n}$ be a trajectory of length $n$. Then if $\varepsilon^{\prime}=\varepsilon\bullet\left(\mathbf{a}:\mathbf{s}\right)$ and $\pi^{\prime}=\pi\bullet\left(\mathbf{s}:\mathbf{a}\right)$ are valid, then $\pi^{\prime}$ is optimal for $\varepsilon^{\prime}$.
Proof.
As previously it suffices to prove the claim for evolution by a single pair $\left(a,s\right)$ and get the general statement by induction.
Given a policy $\pi_{0}\colon T$ by the value-laden Bellman theorem we know that
$S\left(\pi_{0},\varepsilon\bullet\left(a,\left(s,\bot\right)\right)\right)\leq S\left(\pi\bullet\left(s,a\right),\varepsilon\bullet\left(a,\left(s,\bot\right)\right)\right).$ (97)
We can multiply this inequality
$\frac{\mathbb{P}_{\varepsilon}\left(\left(s,\bot\right)\right)}{\mathbb{P}_{\varepsilon}\left(\left(s,\bot\right)\right)+\mathbb{P}_{\varepsilon}\left(\left(s,\top\right)\right)}S\left(\pi_{0},\varepsilon\bullet\left(a,\left(s,\bot\right)\right)\right)\leq\\ \frac{\mathbb{P}_{\varepsilon}\left(\left(s,\bot\right)\right)}{\mathbb{P}_{\varepsilon}\left(\left(s,\bot\right)\right)+\mathbb{P}_{\varepsilon}\left(\left(s,\top\right)\right)}S\left(\pi\bullet\left(s,a\right),\varepsilon\bullet\left(a,\left(s,\bot\right)\right)\right),$ (98)
and apply Lemma 5.2.5 to both sides getting
$S\left(\pi_{0},\varepsilon\bullet\left(a,s\right)\right)\leq S\left(\pi\bullet\left(s,a\right),\varepsilon\bullet\left(a,s\right)\right)$ (99)
finishing the proof. ∎
It might also be worth seeing why we have to be restricted to single success environments in the above theorem, however we will only show this at the end of the last section, when we have the tools to construct a simple example.
6 Specifiability
We will now explore the notion of policies being specified by teleo-environments, where we say that a teleo-environment specifies a policy if the policy is uniquely optimal for the given teleo-environment, (possibly within some constrained class of transducers).
A main theme of the section is that within the set of all transducers, the specifiable transducers are exactly the deterministic transducers. While this result is in keeping with what one might expect from classical decision theory, it is perhaps counter-intuitive from a cognitive science perspective, since it suggests that there is no special detectable feature of those systems whose behaviour are optimal for some sensor-motor task, compared to those that are not, besides merely being deterministic.
We prove this by showing that, for any given deterministic policy, a teleo-environment can be constructed for which that policy is uniquely optimal. The way this environment is constructed is rather simple: the agent must behave in exactly the given way, otherwise it will enter the ‘doom’ environment in which success never occurs. It is of note that while such a simple specifying teleo-environment exists for any deterministic policy, there may be other, much less trivial, environments that also specify the same behaviour.
The equivalence between specifiable transducers and deterministic systems does not hold in every constrained class. For this reason much of the current section is taken up with providing criteria for constrained classes of transducers within which it does happen. In Section 6.5 we discuss an example of when it does not happen, which is a version of the famous ‘absent-minded driver’ problem in decision theory.
We begin with the definition of specification:
Definition 6.0.1 (Teleo-specification).
Let $T\subseteq S\mathrel{\triangleright}A$ be a constrained set of policies. We will say that a teleo-environment $\varepsilon\colon A\mathrel{\triangleright}\left(S\times G\right)$ $T$-teleo-specifies a policy $\pi\colon T$ iff
$\forall_{\pi^{\prime}\colon T}S\left(\pi^{\prime},\varepsilon\right)\geq S\left(\pi,\varepsilon\right)\implies\pi^{\prime}=\pi.$ (100)
In other words $\pi$ is uniquely optimal among all the policies in $T$. As before we will usually omit the ‘$T$-’.
We will investigate what conditions are sufficient and necessary for specifiability of policies, that is whether there is an environment that specifies a given policy within its class.
6.1 Specifying environments
Before we investigate properties of specifiable policies, we will present two necessary conditions for an environment to specify a policy among all policies (i.e $S\mathrel{\triangleright}A$-teleo-specify it). While these conditions are not even close to sufficient, they inform how we approach constructing multiple examples in the following text, so should help the reader to build better intuitions.
In this section we additionally assume that the action space $A$ contains at least two elements. Otherwise there is only one policy, and any environment teleo-specifies it. Due to this trivializing nature this could be the assumption throughout the whole paper, but it turns out all other theorems work without it, even if their statement becomes somewhat vacuous.
Proposition 6.1.1 (Uncertain success).
Let $\varepsilon\in A\mathrel{\triangleright}\left(S\times G\right)$ teleo-specify the policy $\pi\in S\mathrel{\triangleright}A$. Then for any valid evolution $\left(\star:\left(\mathbf{s}:\mathbf{t}:\mathbf{a}\right)\right)\in\left(\left\{\star\right\}\times S\times G\times A\right)^{n}$ of the coupled system $c\left(\pi,\varepsilon\right)$, the evolution $\left(\star:\left(\mathbf{s}:\bot^{n}:\mathbf{a}\right)\right)$ is also valid.
Proof.
Fix an arbitrary order on $A$. Let us define a function
$\zeta\colon\left(\left(S\times A\right)^{*}\times S\mathrel{\triangleright}A\right)\to S\mathrel{\triangleright}A$ (101)
recursively as
$\zeta\left(\left(\mathbf{s}:\mathbf{a}\right),\pi\right)=\left(\mathbb{P}_{\pi},\left(i,o\right)\mapsto\begin{cases}\zeta\left(\left(\mathbf{s}_{>1}:\mathbf{a}_{>1}\right),\mathbb{T}_{\pi}\left(\mathbf{s}_{1},\mathbf{a}_{1}\right)\right)&\textrm{ if }\left(i,o\right)=\left(\mathbf{s}_{1},\mathbf{a}_{1}\right),\\ \mathbb{T}_{\pi}\left(i,o\right)&\textrm{ otherwise}\end{cases}\right),$ (102)
and the recursion base
$\zeta\left(\varepsilon,\pi\right)=\left(a_{0},\left(i,o\right)\mapsto\upsilon\right)$ (103)
where $a_{0}$ is the least element of $A$ such that $\mathbb{P}_{\pi}\left(a_{0}\right)\neq 1$ (we can choose such an element, because $A$ has at least two elements), $\upsilon=\left(U,\left(i,o\right)\mapsto\upsilon\right)$, and $U$ is the uniform distribution on $A$. Crucially $\zeta\left(\varepsilon,\pi\right)\neq\pi$, because the former assigns probability $1$ to $a_{0}$, while the latter does not.
If $\left(\mathbf{s},\mathbf{a}\right)$ is a valid evolution of $\pi$, then $\zeta\left(\left(\mathbf{s}:\mathbf{a}\right),\pi\right)\neq\pi$ – the sequence of transitions which culminates in mapping with the empty string and thus a different transducer, is always within the support of the probability distribution.
To prove the proposition assume to the contrary, that there exists such an evolution, for which the associated evolution without success is not valid. Since $\left(\star:\left(\mathbf{s}:\mathbf{t}:\mathbf{a}\right)\right)$ is a valid evolution for the whole system, $\left(\mathbf{s}:\mathbf{a}\right)$ is a valid evolution for $\pi$. This means that $\pi^{\prime}=\zeta\left(\left(\mathbf{s}:\mathbf{a}\right),\pi\right)\neq\pi$, so it suffices to show that it has no lesser success probability to finish the proof.
Consider equation (72). Since $\zeta$ does not change the probabilities for the first $n$ steps and $S_{n}$ only depends on these steps, $S_{n}\left(\pi,\varepsilon\right)=S_{n}\left(\pi^{\prime},\varepsilon\right)$. In the second sum, the only component influenced by $\zeta$ is the one related to the evolution $\left(\mathbf{s}:\mathbf{a}\right)$, but in this component $F_{c\left(\pi,\varepsilon\right)}\left(\mathbf{s},\mathbf{a}\right)=0$, because that is the probability of exactly the environment trajectory $\left(\mathbf{a}:\left(\mathbf{s}:\bot^{n}\right)\right)$, which is not a valid evolution, so the probability of it has to be 00. Thus $S\left(\pi,\varepsilon\right)=S\left(\pi^{\prime},\varepsilon\right)$, ending the proof. ∎
This proposition can be interpreted as the environment ensuring that the teleo-specified agent cannot ever be certain of its success – intuitively (and in fact also in the proof) in such a situation the agent could behave arbitrarily, contradicting uniqueness. This is clearly a consequence of teleo-specification only caring whether a success occurs at least once.
Proposition 6.1.2 (Everything is possible).
Let $\varepsilon\in A\mathrel{\triangleright}\left(S\times G\right)$ teleo-specify the policy $\pi\in S\mathrel{\triangleright}A$. Then for any valid evolution $\left(\star:\left(\mathbf{s}:\mathbf{t}:\mathbf{a}\right)\right)\in\left(\left\{\star\right\}\times S\times G\times A\right)^{n}$ of the coupled system $c\left(\pi,\varepsilon\right)$, the associated probability $\mathbb{P}_{\varepsilon\bullet\left(\mathbf{a}:\left(\mathbf{s}:\mathbf{t}\right)\right)}\left(s\right)\neq 0$ for any $s\in S$.
Proof.
We proceed almost identically as in the previous proof. Assume that there is such an $s\in S$ that $\mathbb{P}_{\varepsilon\bullet\left(\mathbf{a}:\left(\mathbf{s}:\mathbf{t}\right)\right)}\left(s\right)=0$. Also pick an arbitrary $a\in A$, such that $\mathbb{P}_{\pi\bullet\left(\mathbf{s}:\mathbf{a}\right)}\left(a\right)\neq 0$. Let $\mathbf{s}^{\prime}=\mathbf{s}\cdot s$ and $\mathbf{a}^{\prime}=\mathbf{a}\cdot a$. By the choice of $a$, $\pi^{\prime}=\zeta\left(\left(\mathbf{s}^{\prime}:\mathbf{a}^{\prime}\right),\pi\right)\neq\pi$. The rest of the reasoning is exactly as in the previous proof, only the fact that $F_{c\left(\pi,\varepsilon\right)}\left(\mathbf{s}^{\prime},\mathbf{a}^{\prime}\right)=0$ stems from the probability of $s$ being zero, instead of the probability of the failures being zero. ∎
This proposition can be interpreted as the environment checking the behaviour of the agent in any possible situation – intuitively the agent could behave arbitrarily conditioning on a sensory input that is impossible, making it not unique.
Note that the environment has to behave this way only on the optimal trajectory – whenever an agent behaves suboptimally, the environment is not constrained in this way.
6.2 Specifiable are deterministic
We first show that any policy that can be decomposed into a mixture, cannot be specifiable.
Definition 6.2.1 (Nontrivially decomposable).
We say a constrained policy $\pi\colon T\subseteq S\mathrel{\triangleright}A$ is nontrivially decomposable if there are policies $\pi_{k}\colon T\subseteq S\mathrel{\triangleright}A$ such that $\pi_{k}\neq\pi$ and numbers $\alpha_{k}\geq 0$ such that $\pi=\sum_{k=1}^{\infty}\alpha_{k}\pi_{k}$, and there exists $k_{0}\in\mathbb{N}$, such that $\alpha_{k_{0}}\neq 0$.
Lemma 6.2.2 (Decomposable cannot be specified).
Consider a constrained policy $\pi\colon T\subseteq S\mathrel{\triangleright}A$ that is nontrivially decomposable. Then $\pi$ is not $T$-specifiable.
Proof.
We will show there is another policy with no less of a chance of success than $\pi$. By Lemma 4.2.8 we can write
$S\left(\pi,\varepsilon\right)=\sum_{k=1}^{\infty}\alpha_{k}S\left(\pi_{k},\varepsilon\right).$ (104)
Since the right hand side is a mean of success probabilities, at least one of them is greater or equal to the success probability of $\pi$. It also cannot be equal to $\pi$ by assumption. This contradicts the uniqueness condition for specifiable policies, finishing the proof. ∎
This result means that if we can show that some transducers within a class can be nontrivially decomposed, then they cannot be specifiable within that class. In particular, deterministic transducers obviously cannot be decomposed, which suggests the following theorem.
Theorem 6.2.3 (Specifiable policies are deterministic).
Consider a class of policies $T\subseteq S\mathrel{\triangleright}A$ in which all nondeterministic transducers are nontrivially decomposable. Then only deterministic policies can be $T$-specifiable.
Proof.
Follows immediately from Lemma 6.2.2. ∎
However, identifying which transducers can be nontrivially decomposed is not a simple task in general, so it’s not obvious which classes of transducers this theorem refers to. In particular it’s not a priori obvious that all nondeterministic transducers can be nontrivially decomposed in the total class of transducers. We will now show that this is the case by constructing an explicit decomposition of any nondeterministic transducer (Proposition 6.2.6 below).
Definition 6.2.4 (Pointwise differing transducers).
We say that two transducers $\pi,\pi^{\prime}\colon I\mathrel{\triangleright}O$ differ pointwise if $\pi\neq\pi^{\prime}$, but there is an $n\in\mathbb{N}$ and a trajectory $\left(\mathbf{i},\mathbf{o}\right)\in I^{n}\times O^{n}$ such that for any $\left(\mathbf{j},\mathbf{q}\right)\in I^{*}\times O^{*}\setminus\left\{\left(\mathbf{i},\mathbf{o}\right)\right\}$ that are valid evolutions for both $\pi$ and $\pi^{\prime}$ we have
$\mathbb{P}_{\pi\bullet\left(\mathbf{j}:\mathbf{q}\right)}=\mathbb{P}_{\pi^{\prime}\bullet\left(\mathbf{j}:\mathbf{q}\right)},$ (105)
that is they differ only in the probability returned after the single trajectory, in all other cases they are equal.
Because of the interplay between valid evolutions and associated probabilities, the equality actually works for all valid evolutions of one of the transducers that do not have $\left(\mathbf{i},\mathbf{o}\right)$ as a prefix.
Definition 6.2.5 (Pointwise decomposition).
We say two pointwise differing transducers $\pi^{\prime},\pi^{\prime\prime}\colon I\mathrel{\triangleright}O$ and a number $\alpha\in\left(0,1\right)$ constitute a pointwise decomposition of $\pi=\alpha\pi^{\prime}+\left(1-\alpha\right)\pi^{\prime\prime}$.
An arbitrary nondeterministic transducer is pointwise decomposable within the total class of transducers.
Proposition 6.2.6 (Nondeterministic are pointwise decomposable).
We can construct a decomposition as in Definition 6.2.5 for any nondeterministic transducer $\pi\colon I\mathrel{\triangleright}O\setminus D$.
Proof.
First assume that the probability distribution $\mathbb{P}_{\pi}$ associated with the transducer is not a point distribution. We will prove the general case later.
In this case the probability distribution admits a nontrivial decomposition itself, say of the form
$\mathbb{P}_{\pi}=\alpha p+\left(1-\alpha\right)q.$ (106)
Without loss of generality we assume $p$ and $q$ have the same support as $\mathbb{P}_{\pi}$. Now we can define a pointwise decomposition of the transducer by just using this distribution decomposition and a trivial decomposition on the transition. Thus, the transducers shall be
$\pi^{\prime}=\left(p,\mathbb{T}_{\pi}\right)$ (107)
and
$\pi^{\prime\prime}=\left(q,\mathbb{T}_{\pi}\right).$ (108)
The decomposition then takes the shape of
$\pi=\alpha\pi^{\prime}+\left(1-\alpha\right)\pi^{\prime\prime}.$ (109)
This equality holds on the mix of probabilities because of their definition and on the transition because the transition functions are equal, so the terms with the normalization in the definition cancel out. It’s trivial to check that the transducers differ pointwise for the empty trajectory.
For the general case note that a nondeterministic transducer has to have a valid evolution $\left(\mathbf{i},\mathbf{o}\right)\in\left(I\times O\right)^{n}$, the result of which $\pi_{0}=\pi\bullet\left(\mathbf{i}:\mathbf{o}\right)$ has a nondeterministic associated distribution. By the previous step we know that $\pi_{0}$ pointwise decomposes, so it would be sufficient to prove that if an arbitrary evolution of a transducer pointwise decomposes, then so does the transducer. We will do that coinductively.
By a similar argument as in the proof of Theorem 5.0.2 it is sufficient to prove this for the case of $n=1$ and the rest follows by induction. Let us then consider $\pi_{0}=\pi\bullet\left(i,o\right)$ such that
$\pi_{0}=\alpha\pi_{0}^{\prime}+\left(1-\alpha\right)\pi_{0}^{\prime\prime}$ (110)
is a pointwise decomposition. We construct a decomposition of the original transducer into
$\pi^{\prime}=\left(\mathbb{P}_{\pi},\left(j,q\right)\mapsto\begin{cases}\pi_{0}^{\prime}&\textrm{ if }\left(j,q\right)=\left(i,o\right),\\ \mathbb{T}_{\pi}\left(j,q\right)&\textrm{ otherwise}\end{cases}\right),$ (111)
and
$\pi^{\prime\prime}=\left(\mathbb{P}_{\pi},\left(j,q\right)\mapsto\begin{cases}\pi_{0}^{\prime\prime}&\textrm{ if }\left(j,q\right)=\left(i,o\right),\\ \mathbb{T}_{\pi}\left(j,q\right)&\textrm{ otherwise}\end{cases}\right).$ (112)
The decomposition is then
$\pi=\alpha\pi^{\prime}+\left(1-\alpha\right)\pi^{\prime\prime},$ (113)
which adds up correctly because the associated probabilities are identical, as are all transitions other than by $\left(i,o\right)$, and that last one is
$\begin{split}&\frac{\alpha\mathbb{P}_{\pi}\left(o\right)\mathbb{T}_{\pi^{\prime}}\left(i,o\right)+\left(1-\alpha\right)\mathbb{P}_{\pi}\left(o\right)\mathbb{T}_{\pi^{\prime\prime}}\left(i,o\right)}{\alpha\mathbb{P}_{\pi}\left(o\right)+\left(1-\alpha\right)\mathbb{P}_{\pi}\left(o\right)}\\ &\qquad=\alpha\pi_{0}^{\prime}+\left(1-\alpha\right)\pi_{0}^{\prime\prime}=\pi_{0},\end{split}$ (114)
where the first equality is by canceling out the equal probabilities and applying the transition defined above. This is exactly the transition we would expect.
It is pointwise, because the decomposition of $\pi_{0}$ is pointwise by assumption, and any other associated probabilities introduced by this construction are equal between $\pi^{\prime}$ and $\pi^{\prime\prime}$. ∎
Among the constrained transducer classes we already introduced nondeterministic one-flip transducers are pointwise decomposable (we just need to decompose them arbitrarily at the flip, as the above construction implies), while nondeterministic UFS transducers (including i.i.d. transducers) are not (intuitively, decomposing a maximally complex unifilar $n$ state transducer $\pi\colon\mathrm{UFS}_{n}\setminus\mathrm{UFS}_{n-1}$ would require adding a special memory state for the point of decomposition).
Proposition 6.2.6 together with Theorem 6.2.3 implies that only deterministic transducers can be specifiable among unconstrained transducers.
6.3 Deterministic are specifiable
In this section we show that every deterministic policy is specifiable by explicitly constructing an environment that specifies it. We use a very similar technique as in Example 5.1.3, in which we showed that UFS transducers do not satisfy filtering.
Definition 6.3.1 (Uniform testing environment).
Given a deterministic policy $\pi\in D\subset S\mathrel{\triangleright}A$ we define its uniform testing environment as
$\varepsilon_{\mathrm{test}}\left(\pi\right)=\\ \left(\frac{1}{4}U_{\bot}+\frac{3}{4}U_{\top},\left(a,\left(s,t\right)\right)\mapsto\begin{cases}\varepsilon_{\mathrm{test}}\left(\pi\bullet\left(s,a\right)\right)&\textrm{ if the evolution is valid,}\\ \varepsilon_{0}&\textrm{ otherwise}\end{cases}\right),$ (115)
where $U_{\bot}$ and $U_{\top}$ are the uniform nothing distribution and the uniform success distribution (Definition 5.1.1) and $\varepsilon_{0}$ is the doom environment (Definition 5.1.2). For deterministic policies that evolution is valid iff $a$ is the action the policy takes at this step.
In other words, this environment returns a random state and, as long as the policy behaves like $\pi$, has a $\frac{3}{4}$ chance of returning a success at every step. The exact value of this probability is irrelevant as long as it differs from zero or one; we picked the value $\frac{3}{4}$ to make a future example simpler.
Theorem 6.3.2 (Deterministic policies are specifiable).
Any deterministic policy $\pi\in D\subset S\mathrel{\triangleright}A$ is specifiable.
Proof.
We will show that $\varepsilon_{\mathrm{test}}\left(\pi\right)$ specifies $\pi$. Since this pair has a success chance of $\frac{3}{4}$ at every step we immediately get
$S\left(\pi,\varepsilon_{\mathrm{test}}\left(\pi\right)\right)=1.$ (116)
It suffices to show that any other policy $\pi^{\prime}\colon S\mathrel{\triangleright}A$ has a nonzero chance of failure. We will show this inductively with respect to the shortest evolution after which the associated probabilities of the two policies differ.
If the length is 00, then $\mathbb{P}_{\pi}\neq\mathbb{P}_{\pi^{\prime}}$. Since $\mathbb{P}_{\pi}$ is a point distribution, say focused on $a$, this means that there is an action $a^{\prime}$, such that $\mathbb{P}_{\pi^{\prime}}\left(a^{\prime}\right)$ is not zero. Since the first step results in success with probability $\frac{3}{4}$ and the doom environment never returns success the probability of failure is at least $\frac{1}{4}\mathbb{P}_{\pi^{\prime}}\left(a^{\prime}\right)$, which is not zero.
For the inductive step we can consider the first step $\left(s,a\right)$ of a $n+1$ length evolution leading to differing probabilities. By definition of the uniform test environment the probability of $(s,\bot)$ in the first step is $\frac{1}{4n}$. The probability of $a$ is $1$, since the distributions have to be equal, because $n+1>0$. Since $\pi^{\prime}\bullet\left(s,a\right)$ differs from $\pi\bullet\left(s,a\right)$ after an evolution of length $n$ by the inductive assumption it has a nonzero probability of failure $p$. Thus $\pi^{\prime}$ has a failure probability of at least $\frac{p}{4n}$, which is clearly nonzero. ∎
In particular this means that for the classes of transducers described in the previous section deterministic transducers exactly coincide with specifiable ones.
6.4 Optimality of determinism
Deterministic policies are optimal in another sense – any environment has a (in general not unique) optimal deterministic policy.
To prove that we will need one more relation on bounded sum sequences and a proof that it respects mixtures and success probability respects it.
Definition 6.4.1.
We define the partial order $\geq$ on $\mathrm{BSS}\left(r_{0}\right)$ coinductively by saying that $\left(r,t\right)\geq\left(r^{\prime},t^{\prime}\right)$ iff $\Sigma\left(\left(r,t\right)\right)\geq\Sigma\left(\left(r^{\prime},t^{\prime}\right)\right)$ and $t\geq t^{\prime}$.
Note that this definition is strictly stronger than just the first sequence having a greater sum – any suffix of this sequence also has to have a greater sum than the corresponding suffix of the second sequence. The main motivation in defining the relation this way is the ability to use it in coinductive reasoning.
Lemma 6.4.2.
Let $\left(r_{k},t_{k}\right),\left(r^{\prime}_{k},t^{\prime}_{k}\right)\colon\mathrm{BSS}\left(s\right)$ be bounded sum sequences and $\alpha_{k}\in\left[0,1\right]$ be such that $\sum_{k=0}^{\infty}\alpha_{k}=1$. If $\forall_{k\in\mathbb{N}}\left(r_{k},t_{k}\right)\geq\left(r^{\prime}_{k},t^{\prime}_{k}\right)$ then
$\sum_{k=1}^{\infty}\alpha_{k}\left(r_{k},t_{k}\right)\geq\sum_{k=0}^{\infty}\alpha_{k}\left(r^{\prime}_{k},t^{\prime}_{k}\right).$ (117)
Proof.
To prove that $\sum_{k=1}^{\infty}\alpha_{k}\left(r_{k},t_{k}\right)\geq\sum_{k=0}^{\infty}\alpha_{k}\left(r^{\prime}_{k},t^{\prime}_{k}\right)$ we have to prove firstly that $\Sigma\left(\sum_{k=1}^{\infty}\alpha_{k}\left(r_{k},t_{k}\right)\right)\geq\Sigma\left(\sum_{k=0}^{\infty}\alpha_{k}\left(r^{\prime}_{k},t^{\prime}_{k}\right)\right)$ and secondly that $\sum_{k=0}^{\infty}\alpha_{k}t_{k}\geq\sum_{k=0}^{\infty}\alpha_{k}t^{\prime}_{k}$. We get the first required inequality by Lemma 4.2.6, and the second by the coinductive hypothesis. ∎
Lemma 6.4.3.
The sum of bounded sum sequences respects the order from Definition 6.4.1, that is
$\left(r,t\right)\geq\left(r^{\prime},t^{\prime}\right)\implies\Sigma\left(\left(r,t\right)\right)\geq\Sigma\left(\left(r^{\prime},t^{\prime}\right)\right).$ (118)
Proof.
It suffices to note that the first requirement for inequality of bounded sum sequences is exactly the desired result. ∎
Theorem 6.4.4 (Deterministic are optimal).
Let $\varepsilon\colon A\mathrel{\triangleright}\left(S\times G\right)$ and $\pi\colon S\mathrel{\triangleright}A$ be an environment-policy pair. Then there exists a deterministic policy $\pi^{\prime}\colon D\subset S\mathrel{\triangleright}A$ such that
$S\left(\pi^{\prime},\varepsilon\right)\geq S\left(\pi,\varepsilon\right).$ (119)
Proof.
We will explicitly construct a mapping $C_{\varepsilon}\colon S\mathrel{\triangleright}A\to D$, such that
$\sigma\left(c\left(C_{\varepsilon}\left(\pi\right),\varepsilon\right)\right)\geq\sigma\left(c\left(\pi,\varepsilon\right)\right),$ (120)
so that setting $\pi^{\prime}=C_{\varepsilon}\left(\pi\right)$ and applying Lemma 6.4.3 will finish the proof.
To do that fix a total order on $A$ – we need this only to make one choice unique, so the order does not have to represent anything in particular. Define $\mathrm{oa}\left(\pi,\varepsilon\right)$ (for ‘optimal action’) as the element of
$\mathrm{argmax}\left\{\sum_{s\in S}\mathbb{P}_{\varepsilon}\left(\left(s,\bot\right)\right)\sigma\left(c\left(\pi\bullet\left(s,a\right),\varepsilon\bullet\left(a,\left(s,\bot\right)\right)\right)\right)\right\}$ (121)
that is least with respect to this order. Now define
$C_{\varepsilon}\left(\pi\right)=\left(c_{\varepsilon}\left(\pi\right),C_{\varepsilon}\circ\mathbb{T}_{\pi}\right),$ (122)
where $c_{\varepsilon}\colon S\mathrel{\triangleright}A\to P\left(A\right)$ is defined as
$c_{\varepsilon}\left(\pi\right)=\begin{cases}1&\textrm{ if }a=\mathrm{oa}\left(\pi,\varepsilon\right),\\ 0&\textrm{ otherwise.}\end{cases}$ (123)
Note that in (122), $\mathbb{T}_{\pi}$ denotes a restriction of the transition function, as usual.
It remains to prove that (120) holds. If we write $(r^{\prime},t^{\prime})=\sigma\left(c\left(C_{\varepsilon}\left(\pi\right),\varepsilon\right)\right)$ and $(r,t)=\sigma\left(c\left(\pi,\varepsilon\right)\right)$ this means we have to prove that $\Sigma\left((r^{\prime},t^{\prime})\right)\geq\Sigma\left((r,t)\right)$ and $t^{\prime}\geq t$. Let us start by considering the second inequality. Expanding the definition of the left hand side gives
$t^{\prime}=\sum_{s\in S}\sum_{a\in A}\mathbb{P}_{c\left(C_{\varepsilon}\left(\pi\right),\varepsilon\right)}\left(s,\bot,a\right)\sigma\left(c\left(C_{\varepsilon}\left(\pi\right)\bullet\left(s,a\right),\varepsilon\bullet\left(a,\left(s,\bot\right)\right)\right)\right).$ (124)
Applying the definition of $C_{\varepsilon}$ to the evolution, while setting $a_{0}=\mathrm{oa}\left(\pi,\varepsilon\right)$ for brevity and noting that the only nonzero probability corresponds to it we have
$t^{\prime}=\sum_{s\in S}\mathbb{P}_{c\left(C_{\varepsilon}\left(\pi\right),\varepsilon\right)}\left(s,\bot,a_{0}\right)\sigma\left(c\left(C_{\varepsilon}\left(\pi\bullet\left(s,a_{0}\right)\right),\varepsilon\bullet\left(a_{0},\left(s,\bot\right)\right)\right)\right).$ (125)
We can then use the coinductive hypothesis (120) to obtain
$t^{\prime}\geq\sum_{s\in S}\mathbb{P}_{c\left(C_{\varepsilon}\left(\pi\right),\varepsilon\right)}\left(s,\bot,a_{0}\right)\sigma\left(c\left(\pi\bullet\left(s,a_{0}\right),\varepsilon\bullet\left(a_{0},\left(s,\bot\right)\right)\right)\right)$ (126)
and finally by the choice of $a_{0}$ and Lemma 6.4.2 this is no less than the weighted mean
$t^{\prime}\geq\sum_{s\in S}\sum_{a\in A}\mathbb{P}_{c\left(\pi,\varepsilon\right)}\left(s,\bot,a\right)S\left(\pi\bullet\left(s,a\right),\varepsilon\bullet\left(a,\left(s,\bot\right)\right)\right)=t,$ (127)
as required. After applying Lemma 6.4.3 this also gives us
$\Sigma\left(t^{\prime}\right)\geq\Sigma\left(t\right)$ (128)
For the first inequality we need notice that
$\sum_{s\in S}\sum_{a\in A}\mathbb{P}_{c\left(C_{\varepsilon}\left(\pi\right),\varepsilon\right)}\left(s,\bot,a\right)=\sum_{s\in S}\sum_{a\in A}\mathbb{P}_{c\left(\pi,\varepsilon\right)}\left(s,\bot,a\right),$ (129)
because the total probability assigned to a state does not depend on the policy. Since $\Sigma\left(\left(r,t\right)\right)=r+\Sigma\left(t\right)$, it remains to use (128) to finish the proof. ∎
As a simple corollary we get.
Proposition 6.4.5 (Deterministic are uniquely optimal if present).
Let $T\subseteq S\mathrel{\triangleright}A$ be a constrained class of transducers such that $D\subseteq T$. Then deterministic transducers are exactly $T$-teleo-specifiable.
Proof.
By Theorem 6.3.2 they are specifiable, and any nondeterministic transducer cannot be uniquely optimal due to Theorem 6.4.4. ∎
This proposition could have been used to prove that deterministic policies are exactly specifiable policies among unconstrained policies instead of Theorem 6.2.3. It works in any class of policies that contains all deterministic transducers, even if the nondeterministic transducers are not decomposable in that class. However, Theorem 6.2.3 tells us more about classes which don’t necessarily contain all deterministic transducers. For example consider one-flip transducers without deterministic transducers – Proposition 6.4.5 tells us nothing about specifiability in this class, while Theorem 6.2.3 clearly shows that there are no specifiable policies there at all.
6.5 The absent-minded driver
If we constrain policies to a set in which nondeterministic policies do not admit nontrivial decompositions, we might encounter policies that are specifiable, but not deterministic. In this section we translate one classic example of such a situation into the language of transducers – the absent-minded driver problem, first proposed in [32]. Our version differs somewhat from the original formulation to make it more suited for our setting and make our analysis simpler, but the reasons underlying the nondeterminism of the optimal solution remain the same.
In plain language the problem can be stated as follows. There is a motorway with infinitely many identical exits. At every exit the driver can either continue or turn. The driver’s goal is to turn at the second exit, but unfortunately they are absent-minded, in the sense that they have no memory at all and hence do not know whether they have already passed any exits or how many. What is the optimal policy for such a driver?
We will need one more generic environment for our example.
Definition 6.5.1 (Success environment).
The success environment is one that returns one success and then becomes the doom environment
$\varepsilon_{1}=\left(U_{\top},\left(a,\left(s,t\right)\right)\mapsto\varepsilon_{0}\right).$ (130)
Note that since we only care about success occuring once, the doom environment after transition is not particularly gloomy.
Example 6.5.2 (The absent-minded driver).
Let $A=\left\{c,e\right\}$ (where $c$ stands for ‘continue’ and $e$ for ‘exit’), and $S=\left\{0\right\}$. We define the ‘please exit’ environment as
$\varepsilon_{e}=\left(U_{\bot},\left(a,\left(s,t\right)\right)\mapsto\begin{cases}\varepsilon_{1}&\textrm{ if }a=e,\\ \varepsilon_{0}&\textrm{ otherwise}\end{cases}\right).$ (131)
The absent-minded driver environment is then
$\varepsilon=\left(U_{\bot},\left(a,\left(s,t\right)\right)\mapsto\begin{cases}\varepsilon_{e}&\textrm{ if }a=c,\\ \varepsilon_{0}&\textrm{ otherwise}\end{cases}\right).$ (132)
Making the absent-minded driver actually absent-minded means we restrict our policies to only constant policies. As mentioned before, they are the same policies as unifilar one state policies $\mathrm{UFS}_{1}$, and the absent-mindedness translates to only having one memory state.
A constant policy is uniquely defined by the probabilities it assigns to actions, let us denote them by $p_{c}$ and $p_{e}$ respectively. A simple computation tells us that the success probability of such a constant policy on $\varepsilon$ is exactly $p_{c}p_{e}$, since the only way of achieving success is to first continue and then exit. The unique values that maximize this probability are $p_{c}=p_{e}=\frac{1}{2}$. Thus the constant policy assigning these probabilities is $\mathrm{UFS}_{1}$-teleo-specified by $\varepsilon$, and it is nondeterministic as promised.
It is worth noting an important point about this example. Since the transducer with $p_{c}=p_{e}=\frac{1}{2}$ is optimal within the constrained class $\mathrm{UFS}_{1}$ for the absent-minded driver problem, we could attribute the absent-minded driver problem as a normative-epistemic state. However, such a normative-epistemic interpretation lacks a property held by the other examples so far, in that its beliefs do not update consistently according to the value-laden filtering scheme described in Section 5. This lack of temporal consistency with respect to value-laden filtering is a consequence of the fact that $\mathrm{UFS}_{1}$ does not have the value-laden Bellman property.
6.6 No sensorimotor-only filtering in generic environments
Recall that in our setting, the agent’s goal is to achieve success at least once, with subsequent success signals counting for nothing. This leads to the notion of ‘value-laden filtering’, in which the agent always behaves as though it has not already achieved success, since this maximises its overall chances of eventually achieving success.
In light of this we shouldn’t expect the evolution of an optimal agent to remain optimal for the environment evolved only by the sensorimotor signals, as that environment becomes a mixture of the evolution after a failure (which the optimal agent cares about) and after a success (which the optimal agent shouldn’t care about). If we, then, construct an environment which requires very different behaviour to achieve success after at least one success than otherwise, we should be able to exhibit a counterexample to sensorimotor-only filtering.
Indeed, as promised at the end of the filtering section, we present such an example based on this approach. First let us formally define a special environment behaving as described above.
Definition 6.6.1 (Tricky testing environment).
Given two deterministic policies $\pi,\pi^{\prime}\in D\subset S\mathrel{\triangleright}A$ we define their tricky testing environment as
$\varepsilon^{\prime}_{\mathrm{test}}\left(\pi,\pi^{\prime}\right)=\\ \left(\frac{1}{4}U_{\bot}+\frac{3}{4}U_{\top},\left(a,\left(s,t\right)\right)\mapsto\begin{cases}\varepsilon^{\prime}_{\mathrm{test}}\left(\pi\bullet\left(s,a\right),\pi^{\prime}\right)&\textrm{ if }t=\bot,\\ \varepsilon_{\mathrm{test}}\left(\pi^{\prime}\right)&\textrm{ if }t=\top,\\ \varepsilon_{0}&\textrm{ if }\pi\bullet\left(s,a\right)\textrm{ is not valid}\end{cases}\right),$ (133)
where $U_{\bot}$ and $U_{\top}$ are the uniform nothing distribution and the uniform success distribution (Definition 5.1.1), $\varepsilon_{\mathrm{test}}\left(\pi^{\prime}\right)$ is the testing environment for $\pi^{\prime}$ (Definition 6.3.1) and $\varepsilon_{0}$ is the doom environment (Definition 5.1.2).
The tricky testing environment $\varepsilon^{\prime}_{\mathrm{test}}\left(\pi,\pi^{\prime}\right)$ behaves like the uniform testing environment for $\pi$ until it produces a success, after which it starts behaving like the testing environment for $\pi^{\prime}$.
It’s easy to see that a tricky testing environment is equal to the related uniform testing environment after truncation, that is $Z\left(\varepsilon_{\mathrm{test}}\left(\pi\right)\right)=Z\left(\varepsilon^{\prime}_{\mathrm{test}}\left(\pi,\pi^{\prime}\right)\right)$. In particular, by Lemma 5.2.4 this means that $\varepsilon^{\prime}_{\mathrm{test}}\left(\pi,\pi^{\prime}\right)$ teleo-specifies $\pi$.
Example 6.6.2 (No sensorimotor-only filtering in an arbitrary environment).
Let $A=\left\{0,1\right\}$, and $\pi,\pi^{\prime}\colon D\subset S\mathrel{\triangleright}A$ be the policies constantly returning 00 and $1$ respectively. Consider then the environment $\varepsilon^{\prime}_{\mathrm{test}}\left(\pi,\pi^{\prime}\right)$. As mentioned above it teleo-specifies $\pi$, in particular this policy is optimal for the environment. However the policy $\pi\bullet\left(s,0\right)=\pi$ is not optimal for
$\varepsilon^{\prime}_{\mathrm{test}}\left(\pi,\pi^{\prime}\right)\bullet\left(0,s\right)=\frac{1}{4}\varepsilon^{\prime}_{\mathrm{test}}\left(\pi\bullet\left(s,0\right),\pi^{\prime}\right)+\frac{3}{4}\varepsilon_{\mathrm{test}}\left(\pi^{\prime}\right).$ (134)
To see this it suffices to compute the success probabilities of $\pi$ and $\pi^{\prime}$ on these environments. Since the policies never return the same value, then by construction of the environments
$S\left(\pi,\varepsilon_{\mathrm{test}}\left(\pi^{\prime}\right)\right)=S\left(\pi^{\prime},\varepsilon^{\prime}_{\mathrm{test}}\left(\pi,\pi^{\prime}\right)\right)=\frac{3}{4},$ (135)
and by the proof of Theorem 6.3.2 combined with Lemma 5.2.4
$S\left(\pi^{\prime},\varepsilon_{\mathrm{test}}\left(\pi^{\prime}\right)\right)=S\left(\pi,\varepsilon^{\prime}_{\mathrm{test}}\left(\pi,\pi^{\prime}\right)\right)=1.$ (136)
Thus, we can finish the proof using Lemma 4.2.8 twice
$S\left(\pi,\varepsilon^{\prime}_{\mathrm{test}}\left(\pi,\pi^{\prime}\right)\bullet\left(0,s\right)\right)=\frac{13}{16}<\frac{15}{16}=S\left(\pi^{\prime},\varepsilon^{\prime}_{\mathrm{test}}\left(\pi,\pi^{\prime}\right)\bullet\left(0,s\right)\right).$ (137)
7 Limitations and future work
In this paper we outlined a formalism for thinking about several ways of describing systems as potential agents and have proven some key results. In this final section we discuss some of the limitations and possible extensions to the work, before concluding.
One can envisage several ways in which the work could be extended at the mathematical level. These include broadening or generalising the notion of norm (for example, in order to consider reward maximisation rather than success probability maximisation, or to consider multiple goals), as well as extensions to continuous time and continuous-valued inputs and outputs. Of particular interest would be considerations of multiple interacting systems, rather than only a single agent and its environment. One could also consider the relationship between the agent’s beliefs about its environment (as attributed to it by the theorist), which we consider here, and the theorist’s beliefs about the same environment, which might be different.
In terms of limitations, one crucial missing component of this work is Dennett’s idea that for some systems — the co-called intentional systems — the intentional stance is much more productive than for others, and in particular more productive than the physical stance. Our current framework only says which teleo-environments can be consistently attributed as beliefs; it doesn’t currently offer any way to say which belief attributions are ‘better,’ or more natural or pragmatic for a given agent’s behaviour. After all, in Theorem 6.3.2 we showed that one can always attribute a set of beliefs that say, in effect, that the agent believes it must act in exactly the way it does act, otherwise it will fail its goal; but we suspect it will rarely if ever be fruitful to explain a real agent’s behaviour in terms of such a belief. We suspect the answer might lie in bounded rationality, which we have only made a start on considering here.
Another, related, limitation of our approach is that we do not model so-called hyperintensional dimensions of goals or beliefs. For instance, it is possible to believe that there are seven biscuits in a jar without believing that the square of the number of biscuits in the jar is 49, although these propositions both refer to the same state of the world. Roughly speaking, in our framework, when we attribute beliefs (i.e. a teleo-environment) to an agent, we are making a claim that the agent behaves consistently with those beliefs, but we are not making any claim that the agent thinks or reasons about those beliefs. The things we call beliefs and goals in humans come equipped with a particular form which is under-determined by the facts that they map to. This notion of representational form, guise, or mode of presentation is related to the concept of information processing, which seems important if one wants to account for patterns of delays and errors in an agent’s behaviour. We suspect that addressing this issue would require significant additional formal machinery.
To conclude, we have given a simple framework in which some version of the intentional stance makes sense. In particular we have shown that, within this framework, an agent that is optimal for some goal (i.e. teleo-environment) will in future time steps be optimal for an updated version of the same goal, where the updating is given by ‘value-laden (Bayesian) filtering.’ We have also shown that an agent is uniquely optimal for some teleo-environment if and only if it is deterministic. However, there is much work to be done in extending the framework and in relating it more clearly to Dennett’s ideas.
ACKNOWLEDGEMENT
This article was produced with financial and technical support from Principles of Intelligent Behaviour in Biological and Social Systems (PIBBSS), and Simon McGregor’s work was supported with a scholarship grant from the Alignment of Complex Systems Research Group (ACS) at Charles University in Prague. Nathaniel Virgo’s work on this publication was made possible through the support of Grant 62229 from the John Templeton Foundation. The opinions expressed in this publication are those of the author(s) and do not necessarily reflect the views of the John Templeton Foundation.
The authors would also like to thank Martin Biehl for feedback on the manuscript and Robert Obryk for fruitful discussions about technical details.